A Bibliography of Publications of Yousef Saad

Yousef Saad

Computer Science Dept., University of Minnesota 4-192 EE/CSci Building, 200 Union Street S.E.

Minneapolis, MN 55455 USA

Tel: +1 612 624 7804 FAX: +1 612 625 0572

E-mail: saad@cs.umn.edu (Internet)

09 July 2025 Version 2.80

Abstract

This bibliography records publications of Yousef Saad.

Title word cross-reference

3D [GHS10]. exp(−τA)b [SSS10]. f(A)b [CAS11]. ILU [LSC03]. ILUS [CS97c]. k [CrFS09]. LU [CS97c, LSS03b, Saa94d]. tr(f(A)) [CS18, UCS17].

’02 [AGPS03].

1982 [KR83]. 1988 [BTS^{+}89]. 1992 [RRV93]. 1993 [BCEP94].

20th [Sv00, BW01].

5 [WS93].

Abaffy [Saa92h]. ABS [Saa92h]. Abstract [SS85c]. accelerated [LS13b]. accelerating [KKPS18]. Acceleration [BRZS18, KS87, Saa84b, Saa25, BCRZS22, CS99, rFS09, HTZ^{+}24, KS92, ZSTC06a]. acceptors [SKBS88]. acoustic [EGMS20]. Adapted [FSUS20]. ADI [MS92, MS93]. advances [GGL94]. algebra [DS91a, RRV93]. Algebraic [LS17, GHS10, LSS03a, SS02b, SST04, SSC04, XLS16]. Algorithm [DS91b, LXV^{+}16, Saa85a, SYEG00, XAKS23, ZS07, ESS86, GS87, GS88b, GS88a, GS89b, Ruh94, Saa74c, Saa80a, Saa82a, Saa86c, SS86c, SL86, SL88, SW93, Saa93a, SW96b, XAKS24, Saa91a]. algorithmes [Saa74b]. Algorithms [AGPS03, ASSS11, BDG^{+}10, CS92, CS85a, CS86, CTJ^{+}95, CTSZ07, CZC^{+}09, LXES19, SS85g, Saa92a, Saa92h, Saa94a, Saa94b, Saa06, BGSS14, BS94, CS93, CS96, FRSY96, GS94, KS87, Saa90b, Saa94e, US19, VS14]. Alternating [JSS87, SS85c]. amplitude [WGSC18]. analogue [CJ04]. Analysis [BSS09, BSS10, BW01, Saa92b, Saa94b, Saa97, Saa16, BJR^{+}09, Saa94e, Saa00b]. analytics [KMB^{+}18]. Anderson [BRZS18, BCRZS22]. Anderson-type [BCRZS22]. angle [LSS86, SL86, SL88]. Application [CS12, CTWS94]. Applications [AGPS03, ASSS11, BKS08, BDG^{+}10, FSUS20, Saa06, SrFS08, BJR^{+}09, CSS02, CCS10, CS98a, CS85b, Saa83a, Saa90b, Saa90d, SAD^{+}00, SS11, SSC04]. approach [GS90a, SLX^{+}22]. approaches [KKPS18]. Approximate [BS02b, BS02c, CS94, CS97d, CS98b, Saa03a, BS02a, CrFS09, CS97f, USS17a]. Approximating [LSY16]. Approximation [CS09b, FSUS20, GS92a, XAKS23, BSS09, CCE^{+}18, CS97a, CS08, EGMS20, GS90b, GS90a, GS92b, GSS03, ITS07, Saa84a, Saa86b, Saa86e, SS11, UMS17, XAKS24]. Approximations [CAS11, Saa92b, GHS10, US19]. Architectures [IS85, IS86b, IS86a, SS86b, GS89d, SS89b]. arising [Saa84a, Saa86b, Saa86e, SMSW00]. ARMS [SS02b, SST04]. Arnoldi [BSS10, DS91b, Saa80c, SSW98]. array [SSS85]. Assignment [DS91b, Saa88d]. Associated [DS91b]. Atom [TZA^{+}06]. Augmented [Saa97, CS97b]. Automated [KXS18]. automatic [GS94, Saa92a].

Banded [SS85e, SS87]. Based [BS05b, HS06, KS07, SZ99b, SrFS08, HTZ^{+}24, JSS07, LXS16, MOKS12, SW93, SW96b]. Basic [PSWF93, Saa90a]. basis [CTS93, CTS94]. Benchmark [SW88b, SW88a, SW90]. Beresford [Saa83c]. between [BS02c]. Beyond [KXS18]. BILUM [SZ99a]. BILUTM [SZ99b]. biorthogonalization [Saa80a, Saa82a]. bisection [CrFS09]. Block [LS03, LSS03b, MS93, SS80, SZ99a, SZ99b, Saa03a, ZS08, CS97d, GS87, GS88b, GS88a, GS89b, Saa80b, SZ01, Saa23, MS92]. Block-ADI [MS93, MS92]. block-partitioned [CS97d]. Boeing [SW89]. Book [Cha96, Saa83c, Saa95]. bordered [CS85b]. Bounds [Saa94b, Saa94e]. brief [Saa20]. Brownian [ACSS12]. Bulk [TZA^{+}06].

calculation [ZSTC06b]. Calculations [ÖBSC03, SCS10, AJT^{+}07, CTS93, CTS94, JKSC99, SSC^{+}96, ZSTC06a]. Carolina [BCEP94]. Centenary [BCEP94]. century [Sv00, BW01]. CFD [CSW00, SST04]. CG [CLS24]. Chain [PSS92, Saa91c]. chains [BGB^{+}10, RGSB08]. charge [BSTC05]. charging [RGSB08]. Chebyshev [ESS86, Saa84b, XAKS23, XAKS24, ZSTC06a, ZSTC06b, ZS07, ZCS14]. Chebyshev-filtered [ZSTC06a, ZSTC06b, ZCS14]. class [HTZ^{+}24]. classes [rFS09]. clusters [CTJ^{+}95, JTD^{+}94]. CM [PSWF93, WS93]. CM-5 [WS93, PSWF93]. Coarse [MS07a]. Coarse-Grid [MS07a]. Coarsening [MS07b, CSZ22, OKLS15, US19]. codes [GS81, GS83, JKSC99, UMS17]. Communication [SS85a, Saa85a, SS85d, Saa86c, SS86b, SM95, SS89a, SS89b]. Community [CS12]. Compensation [MOKS12]. Complement [DKXS18, LS05b, SS99a, GHS10, KLS16, LXS16, Saa07, XKL^{+}22, ZXS21]. complement-based [LXS16]. Complements [BS05a]. Complex [PS85, PS87, Ruh94, Saa83a, Saa84a, Saa86b, Saa86e, Saa87c]. complexities [GS89d]. Complexity [ISS84, ISS86, Saa85a, Saa86c]. Component [JSS07]. Component-based [JSS07]. Computation [BS05a, BKS08, RRV93, Saa74a, XLS18, LLCS02, dlGGS^{+}05]. Computational [PS20, SM95, Fit86]. Computations [BTS^{+}89, FWPS92, PSWF93, SW88a, Saa94a, SW88b, SW90, Saa90a]. Computers [FWPS92, SS02a, AS88, AS89]. Computing [BSTC05, CAS11, Saa92e, Saa95, SSS10, TS11, XS16, XAKS23, ACSS12, CS18, CSZ22, PS07, Saa80c, SLX^{+}22, TS12, XAKS24]. concepts [Saa24]. Concurrent [Saa95]. condition [Saa84a, Saa86b, Saa86e]. Conference [BCEP94, Fit86, KR83, RRV93]. Confined [ÖBSC03]. Conjugate [SS85g, SS85f, SS86a, SYEG00, Saa06, HTZ^{+}24, Saa85c]. Conquer [LS13a]. consistent [ZSTC06a, ZSTC06b]. Constructed [BS05b]. construction [CrFS09]. continuation [CS85b]. contour [KKPS18, LXSdH20]. control [DS91a, Saa90d]. Convergence [BS94, Saa80b]. convergent [BS89]. convex [BSS09]. Cornelius [BCEP94]. correcting [UMS17]. Correction [LS17, PS07, ZXS21]. corrections [LXS16, XLS16]. counts [DPS16, NPS16]. coupled [KS92]. coupling [dlGGS^{+}05]. Crout [LSC03, LS05a]. cubic [SKBS88]. Cucheb [AKS17]. cyclic [GS87, GS88b, GS88a, GS89b].

dans [Saa74b]. Data [SS85a, SS85d, SS86b, SS89a, SS89b, Saa94a, SM95, CrFS09, KMB^{+}18, SS14]. Davidson [SSW98, SS98b, ZS07]. December [BCEP94]. Decomposition [CS92, HS06, KXS18, LS17, Saa94a, TS11, CS93, CS96, KKPS18, LXS16, PS07, Saa92a, SSZ98, UMS17]. decoupling [KS87]. Definite [SS80, VSS14]. Deflated [CS97b, SYEG00]. deflation [Saa88d]. Dense [CS12, ISS84, ISS86, KMB^{+}18]. Dense-Linear-System [ISS86]. Densities [XLS18, BSTC05, LSY16, USS17a]. Density [BKS08, BSK^{+}03, RGSB08, SS11, dlGGS^{+}05]. density-functional [RGSB08]. dependent [BSK^{+}03, RGSB08, dlGGS^{+}05]. Design [Saa87b, SW95, SW96a, Saa87a, SMSW00]. Detection [CS12]. Development [Saa22]. Developments [BW01]. Diagonal [SZ99c, Saa05, TS11, BKS07, TS12]. diagonalization [JKSC99, ZCS14]. diatomic [CTWS94]. dictionary [USS17b]. Dielectric [ÖBSC03]. difference [CTS93, CTS94, CTWS94, JTD^{+}94, SSS85]. Differential [CSS85, CSS87, SS81]. Dimension [CS09a, KCS09, KCS11, Saa83b]. dimensional [CrFS09, LSS86, LXSdH20, SS14]. Dimensionality [KS07, NBS10, SrFS08]. Dirac [SS11]. Direct [SS85e, SS87, SW96b]. Direction [SS85c, JSS87]. disjoint [Saa83d]. Distributed [MS94, Saa92e, Saa94a, SM95, SS98a, SS99a, SS99c, Saa07]. Distributions [CS14]. Divide [LS13a]. Domain [CS92, KXS18, KKPS18, LS17, Saa94a, SSZ98, SZ99b, TS11, CS93, CS96, LXS16, PS07, Saa92a]. Domain-Based [SZ99b]. Domain-Decomposition-Type [TS11]. Dominance [Saa05]. d’origine [Saa74b]. DQGMRES [SW93, SW96b]. dual [Saa92d, Saa94d]. Dynamic [SSW98]. dynamics [ACSS12, CJWS96, JTD^{+}94].

E. [Saa92h]. Editorial [Saa00a, BGSS14]. Effective [CS09a]. Efficient [AJT^{+}07, DPS16, GS90b, GS92b, GS92a, NPS16, dlGGS^{+}05, LSS86]. eigendecomposition [SS14]. eigenelements [Saa80c]. Eigenfaces [SrFS08]. Eigenproblems [ZS07, KCS09, KCS11, SGSM15]. Eigensolutions [Saa85b]. Eigenvalue [BSS10, rFS12, IS85, IS86b, LXV^{+}16, PS89, Ruh94, Saa83c, Saa84b, Saa11b, Saa16, SSF93, XLS18, DPS16, EGMS20, KLS16, KKPS18, NPS16, Saa82b, Saa83e, Saa89b, Saa92g, SSC^{+}96, Saa23, SSF95, SS98b, WSS98, ZS08]. Eigenvalues [BS05a, Saa74a, LXES19]. Electronic [JKSC99, SCS10, AJT^{+}07, CTS93, CTS94, CKV^{+}03, CTSZ07, CZC^{+}09, SSC^{+}96]. element [KSS03, KSSG04]. Elimination [Saa85a, Saa86a, Saa96, Saa86c, Saa86d, Saa92c]. Elliptic [CSS85, CSS87, GS87, GS88b, GS88a, GS89b, GS89d, KS92, SS81, SSS85]. Engineering [PS20]. Enhanced [SS99b, SZ01]. Environments [Saa87b, Saa92e, CS99, Saa87a]. equation [KSS03, KSSG04, LSS86, LXSdH20, SL86, SL88, ZCS14]. Equations [CSS85, GS92a, MS92, MS93, BCRZS22, BS87, BS90, BS91, CSS87, ESS86, GS87, GS88b, GS88a, GS89b, GS89c, GS89a, GS90b, GS90a, GS92b, GS81, GS83, PS07, SS81, SSS85, Saa90c, Saa20]. Eric [Saa95]. Error [Saa94b, CS18, Saa94e, UMS17]. estimate [CS18]. Estimation [UCS17, DPS16, NPS16, USS17a]. estimator [BKS07]. Études [Saa74b]. Evolution [TZA^{+}06, CTSZ07]. Evolving [Saa16]. EVSL [LXES19]. Exact [Saa03a]. examples [CLS24]. excited [BGB^{+}10, SKBS88]. Experimental [CS97e]. exploration [Fit86]. Exponential [Saa92b, CS98a]. Extended [SS85c]. Extraction [CS12]. Extreme [rFS12].

F [Saa95]. Face [KS05a]. faces [KS05a]. Factored [BS02b, BS02c, BS02a]. Factorization [HS06, LS05a, Saa92d, Saa94d]. Factorizations [MOKS12, XAKS23, CCS10, XAKS24]. Fast [CrFS09, USS17a, UCS17, VS14, XLS18, GS87, GS88b, GS88a, GS89b, GS89d, US19]. February [GGL94]. feedback [Saa88d]. Fermi [SS11]. few [Saa94b, Saa94e]. field [ZSTC06a, ZSTC06b]. Filtered [BKS08, rFS12, Saa06, AKS17, ZSTC06a, ZSTC06b, ZCS14]. Filtering [KXS18, LXV^{+}16]. Filters [XS16]. Finding [Saa03a]. finite [CTS93, CTS94, CTWS94, JTD^{+}94, KSS03, KSSG04]. finite-difference [CTWS94]. finite-difference-pseudopotential [JTD^{+}94]. first [AJT^{+}07]. first-principles [AJT^{+}07]. Five [Saa24]. fixed [Saa25]. fixed-point [Saa25]. flexible [Saa91a, Saa93a]. flow [WGSC18]. flows [LLCS02]. fMRI [SS14]. forces [CJWS96]. format [CS97c]. free [ZCS14]. frequency [LXSdH20]. Function [XS17, SS11]. Functional [BKS08, BSK^{+}03, RGSB08, SS11, dlGGS^{+}05]. Functions [FSUS20]. Further [BSS10, Saa00b].

Gaussian

[CS14, Saa85a, Saa86c, Saa86a, Saa86d].

General

[CS92, CS94, LSC03, Saa94b, Saa96, SZ99a, SZ99b, SS99a, SS02a, CS93, CS96, Saa92a, Saa92c, Saa94c, Saa94e, SSZ98, SZ99c, SZ01, SS02b, Saa07, XKL^{+}22, ZXS20, ZXS21]. Generalized [XLS18, SS86c]. Globally [BS89]. GMRES [Saa91a, CLS24, SS86c, Saa93a, YXS21]. GPU [AKS17, LS13b]. GPU-accelerated [LS13b]. Gradient [SS85g, SS85f, SS86a, SYEG00, Saa85c]. Gradient-like [SS85g]. Gram [Saa86e]. Graph [CSZ22, FSUS20, HS06, SrFS08, VSS14, CrFS09, GS94, OKLS15]. Graph-Based [SrFS08]. Greedy [MS07b, MS07a]. Grid [MS07a]. Guest [BGSS14].

Hand [Saa87d, KMB^{+}18]. harmonic [CJ04]. Harnessing [BGB^{+}10]. Harwell [SW89]. Havsbad [KR83]. Held [KR83, GGL94]. Helmholtz [KSS03, KSSG04, LXSdH20, OKS10]. Hermitian [LXV^{+}16, Saa74a]. Heuristic [GS94]. Hierarchical [DKXS18, HS06]. hierarchy [CCE^{+}18]. High [CSW00, CrFS09, LXSdH20, SS14]. high-dimensional [SS14]. high-frequency [LXSdH20]. High-order [CSW00]. Higher [CTWS94, SKBS88, JTD^{+}94]. Higher-order [CTWS94, JTD^{+}94]. Highly [Saa94c]. Historical [BW01, Saa20]. Houston [Fit86]. Hybrid [BS87, BS90, ESS86, GHS10]. hydrodynamic [ACSS12]. Hypercube [CS85a, CSS85, CS86, CSS87]. Hypercubes [SS85a, SS85d, SS85b, Saa86a, SS88, Saa86d, SS89a].

III [Ruh94]. ILU [CSW00, CS97e, HS06, LS05a, MS94, OKLS15, Saa92d, Saa92c, Saa96, SZ99a, SZ99c, SZ01, Saa03a, Saa05]. ILUM [Saa92c, Saa96]. ILUs [BS02c, BS05b]. ILUT [Saa92d, Saa94d, SZ99b]. IMA [GGL94]. Impact [IS85, IS86b, IS86a]. Implementation [LXES19, AKS17, BSK^{+}03]. Implementations [SS85f, SS86a, Saa91b, Saa93b]. Implicitly [SSW98]. Improving [USS17b]. incoherence [USS17b]. Incomplete [LS06, MOKS12, CCS10, CS97c, Saa92d, SW93, Saa94d, SW96b]. Incremental [CCS10]. Indefinite [DKXS18, XS17, CS97e, Saa83d, Saa84c, Saa88a, Saa88b, Saa88c]. Indexing [SrFS08, VS14]. industrial [SAD^{+}00]. Inexact [WSS98]. Initio [ÖBSC03, JTD^{+}94]. inner [Saa91a, Saa93a]. inner-outer [Saa91a, Saa93a]. Institute [BTS^{+}89]. integration [KKPS18, LXSdH20]. interactions [ACSS12]. Interior [rFS12]. International [BCEP94]. interval [DPS16, NPS16]. intervals [Saa83d]. Invariant [BKS08, PS07]. Inverse [BS02b, BS05b, CS94, CS98b, TS11, BS02a, CS97d, CS97f, TS12]. Inverse-Based [BS05b]. Inverses [BS02c]. Invert [PS87, PS85]. Iron [TZA^{+}06]. irregularly [FRSY96]. issue [ASSS11, BDG^{+}10]. Iteration [Saa16, ZSTC06b, ZCS14]. Iterations [BKS08, CS98b, Saa00b, Saa25]. Iterative [BTS^{+}89, CS85b, GS81, GS83, SS81, Saa83d, SM95, Sv00, Saa03b, Saa20, CSS02, GGL94, JSS07, KMB^{+}18, LS13b, SW94, SW95, SW96a, SKL^{+}97, Saa01, Saa24, Cha96].

J. [Saa92h]. Jacobi [SS98b, Saa23]. January [Fit86]. journey [Saa20].

Kent [RRV93]. Kernels [SM95]. key

[Saa24]. kit [Saa90a]. Kohn

[SCS12, ZCS14]. Krylov [ACSS12, BSS09, BS87, BS89, BS90, BS94, CS99, CCSY98, CS97b, CS14, ESS86, GS92b, GS92a, Ruh94, Saa81, Saa84c, Saa89a, Saa90b, Saa90d, Saa91b, Saa92b, Saa92e, Saa92f, Saa93b, Saa97, Saa98, Saa11a, Saa22, ZS08].

Laguerre [SSS10]. Lanczos

[BCEP94, AKS17, BGB^{+}10, BSTC05, BKS08, CrFS09, CS09a, CS18, rFS12, LXV^{+}16, RGSB08, Saa80a, Saa80b, Saa82a, Saa87d, Saa94b, Saa94e, UCS17]. Lanczos-Type [Saa94b, Saa94e]. Large [BKS08, BTS^{+}89, DS91b, IS86a, LS06, ÖBSC03, PS89, Saa82b, Saa85b, Saa11b, SSF93, ZS07, DS91a, LSY16, Saa74a, Saa80a, Saa80c, Saa81, Saa82a, Saa83b, Saa83e, Saa89b, Saa90c, Saa92g, SSC^{+}96, SAD^{+}00, SSF95, UMS17, WSS98, ZS08]. Latent [SrFS08, VS14]. learned [USS17b]. learning [CSZ22]. Least [CAS11, LS06, Saa83a, Saa87c, XS16, Saa84a, Saa86b, Saa86e]. Least-Squares [LS06, XS16]. level [SSZ98, SZ99c, SZ01]. Library [LXES19, SW94, SW95, SW96a, SKL^{+}97]. Like [DS91b, SS85g]. Linear [Cha96, DKXS18, ITS07, ISS84, ISS86, MS92, MS93, MS94, SS85g, SS85e, SS87, SS98a, SZ99a, SS99a, SS99c, SS02a, XS17, AS88, DS91a, ESS86, GS81, GS83, GSS03, JSS07, KMB^{+}18, LS13b, OKS10, RRV93, Saa81, Saa83d, Saa84c, SSS85, SS86c, Saa87c, Saa88d, Saa88a, Saa88b, Saa88c, SSZ98, SZ99c, SS99b, Sv00, SZ01, Saa01, SS02b, Saa03b, Saa07, Saa20, Saa24, SMSW00, ZXS20, ZXS21]. liquid [LLCS02]. localized [CJWS96]. Low [CS09b, DKXS18, LS13a, LS17, UMS17, CS08, LXS16, XLS16, XKL^{+}22, ZXS20, ZXS21]. Low-Rank [LS13a, LS17, LXS16, XLS16, XKL^{+}22, ZXS20, ZXS21]. LR [Saa74b]. LU [CCS10]. Lyapunov [Saa90c].

machine [CSZ22]. Magnetism [TZA^{+}06]. March [GGL94, KR83, RRV93]. Markov [PSS92, Saa91c]. Massively [FWPS92]. Material [SÖS^{+}00]. Materials [PS20, SCS10]. mathematical [Fit86, Fit86]. Matrices [CS92, CS94, LSC03, LS13a, ÖBSC03, PS87, Saa85b, SW89, Saa96, SZ99b, Saa16, BSS09, CS93, CS96, CS97d, CS97e, LS05a, LSY16, PS85, Ruh94, Saa74a, Saa80c, Saa84a, Saa86b, Saa86e, Saa92c, Saa94c, UMS17, XLS16, XKL^{+}22]. Matrix [AGPS03, ASSS11, AEKS90, BDG^{+}10, FSUS20, FWPS92, IS86a, OKLS15, PSWF93, SW88a, Saa92b, Saa94a, SW94, TS11, XAKS23, BJR^{+}09, BKS07, BGSS14, CCE^{+}18, CS98a, Saa83a, Saa83b, SW88b, Saa90a, SW95, SW96a, SAD^{+}00, TS12, USS17a, US19, VSS14, XAKS24, dlGGS^{+}05, KR83]. Memory [Saa87b, SM95, Saa87a]. Message [Saa87b, Saa87a, WS93]. Method [SS80, Saa87d, CTS93, CTS94, CTWS94, CS18, EGMS20, JTD^{+}94, KSS03, KSSG04, LSS86, Saa80c, Saa85c, Saa23, SCS12, TS12, ZS08, ZCS14]. Methods [BTS^{+}89, Cha96, CCSY98, CS14, DS91b, GS92a, LS17, PSS92, SS81, SS85c, SS85e, SS85f, SS86a, Saa87b, SS87, Saa91b, Saa92e, Saa93b, Saa97, SCS10, Saa11a, Saa11b, Saa22, SSW98, SÖS^{+}00, TS11, ACSS12, BSS09, BS87, BS89, BS90, BS91, CSS02, CS85b, rFS09, Fit86, GS90b, GS92b, GGL94, JSS87, JSS07, KS92, KCS09, KCS11, Saa80a, Saa80b, Saa81, Saa82a, Saa82b, Saa83d, Saa83b, Saa83e, Saa84c, Saa87a, Saa88d, Saa89a, Saa90b, Saa90d, Saa91c, Saa92g, Saa92f, Saa98, Saa01, Saa03b, Saa20, Saa24, Saa25, SS98b]. minimal [SS86c, SW93, SW96b]. minimum [Saa00b]. Minneapolis [BTS^{+}89, GGL94]. Minnesota [BTS^{+}89, GGL94]. MIQR [LS06]. Modeling [PSS92, Fit86]. models [Saa91c]. modern [CSS02, SSC04]. modes [SLX^{+}22]. Modification [MOKS12]. Modified [CS99, Saa84a, Saa86b]. module [SW94, SW95, SW96a]. Molecular [CJWS96, BGB^{+}10, JTD^{+}94]. molecular-dynamics [JTD^{+}94]. molecules [CTWS94]. moment [Saa84a, Saa86b]. Multi [Saa96, Saa92c, SSZ98, SZ99c, SZ01]. Multi-Elimination [Saa96, Saa92c]. multi-level [SSZ98, SZ99c, SZ01]. Multicolor [ZXS20, SS99b]. Multielimination [SZ99a]. Multigrid [CS85a, CS86]. Multilevel [BS05b, KXS18, LS06, SZ99a, SZ99b, Saa05, SrFS08, LSS03a, OKLS15, SS02b, SST04, SSC04, US19, XLS16, XKL^{+}22]. multiple [KMB^{+}18]. Multiprocessor [CS85a, CSS85, CS86, ISS84, ISS86, CSS87]. Multiprocessors [SS85c, Saa85a, JSS87, SS81, Saa86c]. multisecant [rFS09]. Multistage [HS06]. Multivariate [CS14].

N [Saa83c]. nanocrystals

[CTSZ07, CZC^{+}09]. Neighborhood [KS07, KS05b]. News [Cha96, Saa95]. Newton [BS94, WSS98]. nlTGCR [HTZ^{+}24]. NN [CrFS09]. Non [SS99c, SLX^{+}22]. non-perturbative [SLX^{+}22]. Non-standard [SS99c]. nonlinear [BCRZS22, BS87, BS89, BS90, BS91, BS94, EGMS20, rFS09, HTZ^{+}24, KS92, SGSM15]. Nonsymmetric [LSS03b, MS92, MS93, MS07b, Saa84b, SS85g, Saa85b, ESS86, Ruh94, Saa83a, Saa84c, SS86c, Saa87c, Saa88a, Saa88b, Saa88c, Saa89b]. normal [BSS09, SLX^{+}22]. North [BCEP94]. null [ITS07]. null-space [ITS07]. number [Saa86e]. numbers [Saa84a, Saa86b]. Numerical [BW01, PSS92, RRV93, Saa83b, Saa87b, Saa89b, Saa90c, Saa92g, SCS10, Saa11b, Saa87a, Saa91c].

oblique [Saa80a, Saa82a]. Observer [DS91b]. October [BTS^{+}89]. ODE [GS81, GS83]. Ohio [RRV93]. Operator [Saa92b, CS98a]. OPRA [KS05a]. OPRA-faces [KS05a]. Optimal [CS09b, CS08]. Optimization [NBS10, NBS12, BSS09, KCS09, KCS11]. order [CSW00, CTWS94, JTD^{+}94]. Origin [Saa22, Saa74c]. Orthogonal [CS09b, KS05b, KS07, CS08, Saa83d]. orthogonalization [SW93, SW96b]. other [Saa80a, Saa82a]. outer [Saa91a, Saa93a]. Overlapping [CS92, CS93, CS96, LS05b]. overview [Saa90d].

P SPARSLIB [SW94, SW95, SW96a, SKL^{+}97]. Package [SW88a, SS02a, SW88b, SW90, XKL^{+}22]. papers [GGL94]. Parabolic [GS92a, GS89c, GS89a, GS90b, GS90a, GS92b]. Parallel [BDG^{+}10, BGSS14, BSK^{+}03, CSS02, CS97f, FWPS92, FRSY96, GS90a, HS06, IS85, IS86b, IS86a, SS85e, SS85f, SS86b, SS86a, Saa87b, SS87, SW94, SS99c, Saa01, SS02a, SÖS^{+}00, XAKS23, ZSTC06a, AS88, AS89, CS99, GS87, GS88b, GS88a, GS89b, GS89c, GS89a, GS89d, GHS10, LSS03a, LLCS02, SS80, Saa87a, SS89b, Saa92c, Saa94c, SW95, SW96a, SKL^{+}97, SS99b, SSC04, XKL^{+}22, XAKS24, AGPS03, ASSS11]. parGeMSLR [XKL^{+}22]. Parlett [Saa83c]. pARMS [LSS03a, SS02a]. Partial [CSS85, DS91b, Saa85b, XS16, XAKS23, CSS87, Saa88d, XAKS24]. partially [BSTC05]. Particle [LLCS02]. partitioned [CS97d]. partitioning [GS94, LLCS02, Saa74a, VSS14]. Passing [Saa87b, Saa87a, WS93]. Pencils [KR83, XAKS23, XAKS24]. Performance [WS93]. periodic [AJT^{+}07]. perturbative [SLX^{+}22]. Phase [WGSC18]. physical [CSS02, SSC04]. Pite [KR83]. Pivoting [BS02b, BS02a, LS05a]. plane [JKSC99, Saa83a, Saa84a, Saa86b, Saa86e, Saa87c]. plane-wave [JKSC99]. planets [SLX^{+}22]. PMAA [AGPS03]. PMAA’10 [ASSS11]. Point [LS03, LSS03b, Saa25]. pole [Saa88d]. Polynomial [BKS08, CAS11, FSUS20, LXV^{+}16, YXS21, GS90b, LXSdH20, Saa85c]. polynomials [Saa83d, Saa83a, Saa87c, SSS10]. portable [SKL^{+}97]. Positive [SS80, VSS14]. posteriori [CS18]. potential [CTS93, CTS94]. power [ZXS21]. Practical [BTS^{+}89, Saa84c, Saa85c, BTS^{+}89]. Preconditioned [CCSY98, CS14, SS85f, SS86a, Saa91b, Saa93b, Saa98, LS13b, Saa91a, Saa92f, Saa93a]. Preconditioner [BS02b, DKXS18, LS05b, LS06, Saa96, SZ99a, SZ99b, XS17, BS02a, CS97c, Saa92c, XLS16, ZXS20, ZXS21]. Preconditioners [BS05b, CS94, CS98b, LS13a, LS17, LS03, LSS03b, MS92, MS93, MS94, CS97a, CSW00, CS97e, CS97f, GSS03, LXS16, Saa94c, SZ99c, Saa07]. Preconditioning [CS98a, KSS03, KSSG04, OKS10, Saa88a, Saa88b, Saa88c, SAD^{+}00, Saa03a, SMSW00, SSF93, YXS21, LXSdH20, OKLS15, SS99b, SZ01, SSF95, VSS14, WSS98, XKL^{+}22]. preconditionings [Saa85c]. Predicting [SÖS^{+}00, CTJ^{+}95]. Preserving [CCSY98, KS07, KS05b]. Prewhitening [SS14]. primitives [WS93]. principles [AJT^{+}07]. probing [TS12]. Problem [NBS10, NBS12, CKV^{+}03, Saa23, SCS12, Saa83c]. Problems [BSS10, DS91b, rFS12, GGL94, IS85, LS06, LXV^{+}16, LS03, LSS03b, MS07b, PS89, Saa84b, Saa11b, Saa16, SSF93, XLS18, CSW00, DS91a, EGMS20, FRSY96, IS86b, KLS16, KKPS18, Ruh94, Saa82b, Saa83a, Saa83b, Saa83e, Saa89b, Saa90d, Saa92g, SSC^{+}96, SAD^{+}00, SST04, SSF95, WSS98, ZS08]. Procedure [rFS12, AKS17]. procedures [HTZ^{+}24]. Proceedings [BTS^{+}89, KR83, Fit86, RRV93, BCEP94]. Process [BSS10]. Processing [FSUS20]. processors [SSS85]. Projection [BS91, KS07, Saa82b, Saa83e, Saa88d, Saa91c, Saa92h, ITS07, Saa80a, Saa82a]. Projection-Based [KS07]. Projections [KS07, KS05b]. Properties [SS85b, SS88, SÖS^{+}00, CTJ^{+}95, CTSZ07, CZC^{+}09]. Proxy [YXS21]. Proxy-GMRES [YXS21]. pseudo [CTS93, CTS94]. pseudo-potential [CTS93, CTS94]. pseudopotential [CTWS94, JTD^{+}94]. pseudopotentials [CKV^{+}03]. PSPARSLIB [SS98a]. purpose [Saa92a].

QR [LS06, Saa74b]. Quadrature [UCS17]. quantum [CJWS96]. Quasi [SW93, SW96b]. Quasi-minimal [SW93, SW96b].

Raleigh [BCEP94]. Rank [CS09b, DKXS18, LS13a, LS17, CS08, LXS16, USS17b, UMS17, XLS16, XKL^{+}22, ZXS20, ZXS21]. ranks [USS17a]. rates [Saa80b]. Ratio [NBS10, NBS12]. Rational [GSS03, KXS18, Ruh94, SS11, XS16, XS17, EGMS20, GS90a]. Real [PS87, CKV^{+}03, PS85, Ruh94]. recognition [KS05a]. recursive [CrFS09, LSS03a, SS02b, SST04, SSC04]. recycling [SGSM15]. Reduction [CS09a, KS07, NBS10, SrFS08, GS87, GS88b, GS88a, GS89b, KCS09, KCS11]. Relations [BS02c]. reordering [OKLS15]. Reorderings [Saa05]. reorthogonalized [BSTC05]. reservoir [Fit86]. Residual [Saa06, SS86c, SW93, SW96b, Saa00b]. Residual-type [Saa06]. residuals [HTZ^{+}24]. Restart [LXV^{+}16]. Restarted [SSW98]. Restarting [SSW98, SS98b]. Restricted [LS05b]. results [CJ04]. retrieval [WGSC18]. Review [Saa83c, Saa92h]. Reviews [Cha96, Saa95]. Revisiting [Saa23]. reweighted [WGSC18]. Right [Saa87d, KMB^{+}18]. Right-Hand [Saa87d, KMB^{+}18]. Ring [ISS84, ISS86]. Ritz [CJ04]. Robust [SSF93, SSF95, SZ99c]. rotating [SLX^{+}22]. rotation [Saa23].

Saad [Cha96, CJ04]. Saddle [LS03, LSS03b]. Sampling [CS14, US19]. scalable [KMB^{+}18]. Scale [BTS^{+}89]. Schur [BS05a, DKXS18, GHS10, KLS16, LS05b, LXS16, SS99a, Saa07, XKL^{+}22, ZXS21, ZS08]. SchurRAS [LS05b]. Science [PS20]. Scientific [RRV93, Saa95, CSZ22]. seismic [Fit86, SLX^{+}22]. Selection [MS07a]. Self [ZSTC06b, ZSTC06a]. Self-consistent-field [ZSTC06b, ZSTC06a]. Semantic [SrFS08, VS14]. semiconductor [KS87]. semiconductors [SKBS88]. separation [CCE^{+}18]. Sequence [BRZS18]. sets [SS14]. Several [Saa87d]. Sham [SCS12, ZCS14]. Shanks [BRZS18, BCRZS22]. shaped [Saa24]. Shared [Saa87b, Saa87a]. Shift [PS87, PS85]. Shifts [Saa74c, Ruh94]. shrinkage [USS17b]. Si [JTD^{+}94]. Sides [Saa87d, KMB^{+}18]. Signal [FSUS20]. simulation [KS87]. simulations [ACSS12, JTD^{+}94]. Singular [CS09a]. skyline [CS97c]. Slicing [LXES19, SCS12]. Smallest [BS05a]. SMASH [CCE^{+}18]. SNAP [ITS07]. Software [AEKS90, LXES19, Saa92a]. solid [LLCS02]. solid-liquid [LLCS02]. Solution [DS91a, GS92a, ISS84, IS85, ISS86, IS86b, SSC^{+}96, SS98a, SS99c, GS87, GS88b, GS88a, GS89b, GS89c, GS89a, GS90b, GS90a, GS92b, GS81, GS83, ITS07, KSS03, KSSG04, SS81, Saa83d, Saa83b, Saa89b, Saa90c, Saa91c, SW95, SW96a, Sv00, SST04, Saa24, SGSM15, XKL^{+}22]. solver [KMB^{+}18, LSS03a, SS02b, SSC04]. Solvers [SM95, GS89d, GHS10, KKPS18, LS13b, SW94, SKL^{+}97, SST04]. Solving [AS88, AS89, CSS85, CSS87, LXSdH20, MS92, MS93, PS89, SS80, Saa84b, SS85g, SS85e, SSS85, Saa87d, SS87, SS02a, BS91, CS85b, EGMS20, ESS86, LSS86, Saa80a, Saa81, Saa82a, Saa82b, Saa83a, Saa83e, Saa84c, SS86c, SL86, Saa87c, SL88, ZCS14]. Some [GS89d, SW89, Saa92b, BSS09, Saa84c, Saa86e]. SOR [MS94]. Space [YXS21, CKV^{+}03, ITS07]. SPARK [SW90]. Sparse [AEKS90, CS92, Cha96, CS94, CS98b, FWPS92, GHS10, GGL94, IS86a, LSC03, LS06, MS92, MS93, MS94, PSWF93, PS89, SW88a, SW89, Saa94a, SW94, SM95, Saa96, SS98a, SZ99a, SZ99b, SS99a, SS99c, SS02a, XS17, AS88, AS89, CS93, CS96, CS97c, GSS03, JSS07, LS05a, Saa82b, Saa83a, Saa83e, SW88b, SW90, Saa90a, Saa92c, Saa94c, SW95, SW96a, SKL^{+}97, SSZ98, SZ99c, SAD^{+}00, SZ01, Saa01, SS02b, Saa03b, Saa07, SSF95, XLS16, XKL^{+}22, ZXS20, ZXS21, ZCS14]. Sparse-Sparse [CS98b]. SPARSKIT [Saa90a]. Special [ASSS11, BJR^{+}09, BDG^{+}10]. Spectra [XS16, CJWS96]. Spectral [BS05a, KLS16, SGSM15, XLS18, XAKS23, LSY16, USS17a, XAKS24]. Spectrum [DS91b, FSUS20, SCS12]. Spectrum-Adapted [FSUS20]. Spedicato [Saa92h]. Squares [CAS11, LS06, XS16, Saa83a, Saa84a, Saa86b, Saa86e, Saa87c]. standard [SS99c]. Standards [AEKS90]. state [Saa88d]. states [BGB^{+}10, SKBS88]. Statistics [SW89]. Stewart [CJ04]. Stiefel [SS80]. Stochastic [UCS17]. Strategies [MS07b, MOKS12, PS87, SS99c, LLCS02, PS85, SZ01, SGSM15, SMSW00]. Strategy [MS07a]. structural [CTJ^{+}95]. Structure [SCS10, AJT^{+}07, CTS93, CTS94, CKV^{+}03, JKSC99, SSC^{+}96]. Structured [GGL94, CCE^{+}18, FRSY96]. Structures [Saa94a, SM95, Saa03a]. study [CS97e]. Subgraph [CS12]. Subspace [CCSY98, CS14, Saa91b, Saa92b, Saa92e, Saa93b, Saa97, Saa11a, Saa16, Saa22, ACSS12, BSS09, BS89, CS97b, ESS86, Saa81, Saa84c, Saa89a, Saa90b, Saa90d, Saa92f, Saa98, Saa23, ZSTC06a, ZSTC06b, ZCS14]. Subspaces [BKS08, PS07]. Substructuring [KXS18]. sum [CS97a]. Supercomputer [BTS^{+}89, Saa91b, Saa93b]. Supercomputers [PS89, Saa89a]. SVD [CS08, CS09b]. Sweden [KR83]. Sylvester [DS91b]. Sylvester-Like [DS91b]. Symmetric [LS13a, LSS03b, Saa83c, Saa87d, SSF93, ZS07, KLS16, KKPS18, LS05a, Saa83d, Saa23, SSF95, SS98b, VSS14, WSS98, XLS16, ZS08]. Symmetry [CCSY98]. System [ISS84, ISS86, BS87, ITS07, KMB^{+}18]. Systems [DKXS18, MS92, MS93, MS94, ÖBSC03, SS80, SS85g, SS85e, Saa87d, SS87, SS98a, SZ99a, SS99a, SS99c, SS02a, XS17, AJT^{+}07, AS88, AS89, BCRZS22, BS90, BS91, CS85b, CJWS96, ESS86, GSS03, JSS07, KS92, OKS10, Saa80a, Saa81, Saa82a, Saa83d, Saa84c, SS86c, Saa87c, Saa88a, Saa88b, Saa88c, SSZ98, SZ99c, SS99b, Sv00, SZ01, Saa01, SS02b, Saa03b, Saa07, Saa20, Saa24, SMSW00, VSS14, ZXS20, ZXS21, Cha96].

Technique [KS07]. Techniques

[IS86a, Saa84b, SS99a, BCRZS22, CS97b, CS97d, KLS16, KSS03, KSSG04, KS87, Saa74a, Saa88a, Saa88b, Saa88c, SSZ98, SZ99c, SS99b, SAD^{+}00, SS98b, WSS98]. Tensor [CS97a, CS09b, CS08]. Tensors [CS09b, CS08]. Texas [Fit86]. their [GS89d, Saa87c]. Theoretical [Saa94b, Saa94e]. Theory [BKS08, BSS09, BS94, BSK^{+}03, RGSB08, Saa90b, SS11, dlGGS^{+}05]. thermoacoustics [SGSM15]. Thick [LXV^{+}16, SSW98]. Thick-Restart [LXV^{+}16]. three [LSS86, LXSdH20]. three-dimensional [LSS86, LXSdH20]. Threshold [MOKS12, Saa92d, Saa94d, SZ99c].

Threshold-based [MOKS12]. time [BSK^{+}03, RGSB08, dlGGS^{+}05]. time-dependent [BSK^{+}03, RGSB08, dlGGS^{+}05]. tire [SMSW00]. tool [Saa90a]. Tools [SÖS^{+}00, Saa92a]. Topological [SS85b, SS88]. Trace [KCS09, KCS11, NBS10, NBS12]. Transformations [BRZS18]. translations [Saa74b]. trends [Saa92f]. triangular [AS88, AS89]. Turbo [RGSB08]. Two [rFS09, Saa83d]. Type [Saa94b, TS11, BCRZS22, Saa94e, SSZ98, Saa06].

understanding [CLS24]. Unstructured [MS94]. unsymmetric [Saa80a, Saa80c, Saa81, Saa82a]. updating [VS14]. USA [RRV93]. use [Saa84c, Saa85c, Saa87c]. Using [BKS08, CKV^{+}03, SS98a, SSC04, BS05a, CS18, JTD^{+}94, KS05a, LXSdH20, OKLS15, Saa83d, USS17a, UMS17, VSS14, ZSTC06b].

values [VSS14]. Variations [Saa80c, SST04]. Vectors [CS09a, CJ04]. Velde [Saa95]. Version [LS05b, SYEG00, LSS03a]. Versions [LSC03, SZ99a, LS05a]. versus [CS09a]. via [BSS09, CrFS09, CAS11, CS98b, UCS17, USS17b, WGSC18, XAKS23, XAKS24, YXS21, ZSTC06a]. Vibrational [CJWS96, CZC^{+}09]. volume [BJR^{+}09].

wave [JKSC99, LSS86, SL86, SL88]. wide [LSS86, SL86, SL88]. without [CTS93, CTS94, JKSC99, SS14]. Workshop [BTS^{+}89, GGL94].

Yousef [Cha96].

References

Ando:2012:KSM

[ACSS12] Tadashi Ando, Edmond Chow, Yousef Saad, and Jeffrey Skolnick. Krylov subspace methods for computing hydrodynamic interactions in Brownian dynamics simulations. Journal of Chemical Physics, 137(6):064106, August 14, 2012. CODEN JCPSA6. ISSN 1089-7690.

Ashby:1990:SSM

[AEKS90] S. Ashby, H. Elman, D. Kincaid, and Y. Saad. Standards for sparse matrix software. Summary of session: Standards for Sparse Matrix Software, Copper Mountain Conference on iterative methods, April 1990., 1990.

Arbenz:2003:PMA

[AGPS03] Peter Arbenz, Efstratios Gallopoulos, Bernard Philippe, and Yousef Saad. Parallel Matrix Algorithms and Applications (PMAA ’02). Parallel Computing, 29(9):1117–1119, September 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Alemany:2007:EFP

[AJT^{+}07] M. M. G. Alemany, Manish Jain, Murilo L. Tiago, Yunkai Zhou, Yousef Saad, and James R. Chelikowsky. Efficient first-principles calculations of the electronic structure of periodic systems. Computer Physics Communications, 177(4):339–347, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Aurentz:2017:CGI

[AKS17] Jared L. Aurentz, Vassilis Kalantzis, and Yousef Saad. Cucheb: a GPU implementation of the filtered Lanczos procedure. Computer Physics Communications, 220(??):332– 340, November 2017. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465517301984.

Anderson:1988:SST

[AS88] Edward Charles Anderson and Youcef Saad. Solving sparse triangular linear systems on parallel computers. Technical Report CSRD 794, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, June 6, 1988. i + 29 pp.

Anderson:1989:SST

[AS89] E. C. Anderson and Y. Saad.

Solving sparse triangular systems on parallel computers. International Journal of High Speed Computing, 1:73–96, 1989.

Arbenz:2011:SIP

[ASSS11] Peter Arbenz, Yousef Saad, Ahmed Sameh, and Olaf Schenk. Special issue on Parallel Matrix Algorithms and Applications (PMAA’10). Parallel Computing, 37(12):731–732, December 2011. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0167819111001426.

Brown:1994:PCL

[BCEP94] J. David Brown, Moody T. Chu, Donald C. Ellison, and Robert J. Plemmons, editors. Proceedings of the Cornelius Lanczos International Centenary Conference, Raleigh, North Carolina, December 12–17, 1993, volume 73 of Proceedings in Applied Mathematics. SIAM Press, Philadelphia, PA, USA, 1994. ISBN 0-89871-339-0. LCCN QC19.2 .C67 1993.

Brezinski:2022:SAT

[BCRZS22] Claude Brezinski, Stefano Cipolla, Michela Redivo-Zaglia, and Yousef Saad. Shanks and Anderson-type acceleration techniques for systems of nonlinear equations. IMA Journal of Numerical Analysis, 42(4): 3058–3093, October 2022. CO-DEN IJNADH. ISSN 0272-4979 (print), 1464-3642 (elec-

tronic). URL http://academic.
oup.com/imajna/article/42/
4/3058/6354546.

Bekas:2010:SIP

[BDG^{+}10] Costas Bekas, Pasqua D’Ambra, Ananth Grama, Yousef Saad, and Petko Yanev. Special issue on parallel matrix algorithms and applications. Parallel Computing, 36(5–6):213–214, June 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Baroni:2010:HME

[BGB^{+}10] Stefano Baroni, Ralph Gebauer, O. Baris Malcioglu, Yousef Saad, Paolo Umari, and Jiawei Xian. Harnessing molecular excited states with Lanczos chains. Journal of Physics: Condensed Matter, 22(7):074204, February 24, 2010. CODEN JCOMEL. ISSN 1361-648X.

Bekas:2014:PMA

[BGSS14] Costas Bekas, Ananth Grama, Yousef Saad, and Olaf Schenk. Parallel matrix algorithms [Guest editorial]. Parallel Computing, 40(7):159–160, July 2014. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0167819114000647.

Beckermann:2008:SVM

[BJR^{+}09] Bernhard Beckermann, Khalide Jbilou, Lothar Reichel, Yousef Saad, Miloude Sadkane, and Ahmed Salam. Special volume on matrix analysis and applications. Electronic Transactions on Numerical Analysis, 33: vii, 2008/2009. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic).

Bekas:2007:EDM

[BKS07] C. Bekas, E. Kokiopoulou, and Y. Saad. An estimator for the diagonal of a matrix. Applied Numerical Mathematics: Transactions of IMACS, 57(11–12):1214– 1229, November/December 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bekas:2008:CLI

[BKS08] C. Bekas, E. Kokiopoulou, and Yousef Saad. Computation of large invariant subspaces using polynomial filtered Lanczos iterations with applications in density functional theory. SIAM Journal on Matrix Analysis and Applications, 30(1):397– 418, ???? 2008. CODEN SJ-MAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Brezinski:2018:SST

[BRZS18] Claude Brezinski, Michela Redivo-Zaglia, and Yousef Saad. Shanks sequence transformations and Anderson acceleration. SIAM Review, 60(3):646–669, ???? 2018. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brown:1987:HKM

[BS87] Peter N. Brown and Youcef Saad. Hybrid Krylov methods for nonlinear system of equations. Technical Report CSRD 699, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, November 1987. 42 + 4 pp.

Brown:1989:GCN

[BS89] P. N. Brown and Y. Saad. Globally convergent nonlinear Krylov subspace methods. Technical Report 89-57, Research Institute for Advanced Computer Science, NASA Ames, Moffett field, CA, 1989.

Brown:1990:HKM

[BS90] Peter N. Brown and Youcef Saad. Hybrid Krylov methods for nonlinear systems of equations. SIAM Journal on Scientific and Statistical Computing, 11(3):450–481, May 1990. CO-DEN SIJCD4. ISSN 0196-5204.

Brown:1991:PMS

[BS91] Peter N. Brown and Youcef Saad.

Projection methods for solving nonlinear systems of equations. In Nematics (Orsay, 1990), volume 332 of NATO Adv. Sci. Inst. Ser. C Math. Phys. Sci., pages 341–355. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1991.

Brown:1994:CTN

[BS94] Peter N. Brown and Youcef Saad.

Convergence theory of nonlinear Newton–Krylov algorithms. SIAM Journal on Optimization, 4(2):297–330, May 1994. CO-DEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic).

Bollhofer:2001:FAI

[BS02a] Matthias Bollhöfer and Yousef Saad. A factored approximate inverse preconditioner with pivoting. SIAM Journal on Matrix Analysis and Applications, 23(3):692–705, 2001/2002. CO-DEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Bollhofer:2002:FAI

[BS02b] Matthias Bollhöfer and Yousef Saad. A factored approximate inverse preconditioner with pivoting. SIAM Journal on Matrix Analysis and Applications, 23(3):692–705, 2002. CO-DEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/37212.

Bollhofer:2002:RBI

[BS02c] Matthias Bollhöfer and Yousef Saad. On the relations between ILUs and factored approximate inverses. SIAM Journal on Matrix Analysis and Applications, 24(1):219–237, 2002. CO-DEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/37211.

Bekas:2005:CSE

[BS05a] Constantine Bekas and Yousef Saad. Computation of smallest eigenvalues using spectral Schur complements. SIAM Journal on Scientific Computing, 27(2): 458–481, March 2005. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/60352.

Bollhofer:2005:MPC

[BS05b] Matthias Bollhöfer and Yousef Saad. Multilevel preconditioners constructed from inverse-based ILUs. SIAM Journal on Scientific Computing, 27(5):1627– 1650, September 2005. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

URL http://epubs.siam.org/
volume-27/art_60837.html.

Burdick:2003:PIT

[BSK^{+}03] W. Russell Burdick, Yousef Saad, Leeor Kronik, Igor Vasiliev, Manish Jain, and James R. Chelikowsky. Parallel implementation of time-dependent density functional theory. Computer Physics Communications, 156 (1):22–42, December 1–15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465503004132.

Bellalij:2008:ASK

[BSS09] M. Bellalij, Y. Saad, and H. Sadok. Analysis of some Krylov subspace methods for normal matrices via approximation theory and convex optimization. Electronic Transactions on Numerical Analysis, 33:17– 30, 2008/2009. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:

//etna.mcs.kent.edu/vol.33.
2008-2009/pp17-30.dir/pp17-
30.pdf.

Bellalij:2010:FAA

[BSS10] M. Bellalij, Y. Saad, and H. Sadok. Further analysis of the Arnoldi process for eigenvalue problems. SIAM Journal on Numerical Analysis, 48(2): 393–407, ???? 2010. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bekas:2005:CCD

[BSTC05] Constantine Bekas, Yousef Saad, Murilo L. Tiago, and James R. Chelikowsky. Computing charge densities with partially reorthogonalized Lanczos. Computer Physics Communications, 171 (3):175–186, October 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0010465505003103.

Boley:1989:PIM

[BTS^{+}89] Daniel L. Boley, Donald G. Truhlar, Youcef Saad, Robert E. Wyatt, and Lee A. Collins, editors. Practical Iterative Methods for Large Scale Computations: Proceedings of the Minnesota Supercomputer Institute Workshop on Practical Iterative Methods for Large Scale Computations, Minneapolis, 23–25 October 1988. North-Holland Publishing Co., Amsterdam, The Netherlands, 1989. ISBN 0-444-88023-2. LCCN QA297.8.M56 1988. Reprinted from Computer Physics Communications, 53(1– 3), 1989.

Brezinski:2001:NAH

[BW01] Claude Brezinski and Luc Wuytack, editors. Numerical Analysis: Historical Developments in the 20th Century. Elsevier, Amsterdam, The Netherlands, 2001. ISBN 0-444-50617-9 (hardcover), 0-444-59858-8 (e-book). v + 505 pp. LCCN QA297 .N843 2001. URL https://www.

sciencedirect.com/science/
article/pii/B9780444506177500033.

Chen:2011:CLS

[CAS11] Jie Chen, Mihai Anitescu, and Yousef Saad. Computing f(A)b via least squares polynomial approximations. SIAM Journal on Scientific Computing, 33(1): 195–222, ???? 2011. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

URL http://epubs.siam.org/
sisc/resource/1/sjoce3/v33/
i1/p195_s1.

Cai:2018:SSM

[CCE^{+}18] Difeng Cai, Edmond Chow, Lucas Erlandson, Yousef Saad, and Yuanzhe Xi. SMASH: structured matrix approximation by separation and hierarchy. Numerical Linear Algebra with Applications, 25(6):e2204, 25, 2018. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Calgaro:2010:IIL

[CCS10] Caterina Calgaro, Jean-Paul Chehab, and Yousef Saad. Incremental incomplete LU factorizations with applications. Numerical Linear Algebra with Applications, 17(5):811–837, October 2010. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chan:1998:PSP

[CCSY98] T. F. Chan, E. Chow, Y. Saad, and M. C. Yeung. Preserving symmetry in preconditioned Krylov subspace methods. SIAM Journal on Scientific Computing, 20(2):568–581, March 1998. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/31155.

Chan:1996:BNR

[Cha96] Tony F. Chan. Book news & reviews: Iterative Methods for Sparse Linear Systems by Yousef Saad. IEEE Computational Science & Engineering, 3(4):87–90, Winter 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Chen:2004:ARS

[CJ04] Guizhi Chen and Zhongxiao Jia. An analogue of the results of Saad and Stewart for harmonic Ritz vectors. Journal of Computational and Applied Mathematics, 167(2):493–498, June 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S0377042703009403.

Chelikowsky:1996:MDQ

[CJWS96] James R. Chelikowsky, Xiaodun Jing, K. Wu, and Y. Saad. Molecular dynamics with quantum forces: Vibrational spectra of localized systems. Physical Review B: Condensed Matter and Materials Physics, 53(18):12071– 12079, ???? 1996. CODEN PRB-MDO. ISSN 1098-0121. URL

http://link.aps.org/doi/10.
1103/PhysRevB.53.12071.

Chelikowsky:2003:URS

[CKV^{+}03] James R. Chelikowsky, Leeor Kronik, Igor Vasiliev, Manish Jain, and Yousef Saad. Using real space pseudopotentials for the electronic structure problem. In Handbook of numerical analysis, Vol. X, Handb. Numer. Anal., X, pages 613–637. North-Holland Publishing Co., Amsterdam, The Netherlands, 2003.

Carson:2024:TUC

[CLS24] Erin Carson, Jörg Liesen, and Zdeněk Strakoš. Towards understanding CG and GMRES through examples. Linear Algebra and its Applications, 692 (??):241–291, July 1, 2024. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0024379524001381.

Chen:2009:FAG

[CrFS09] Jie Chen, Haw ren Fang, and Yousef Saad. Fast approximate kNN graph construction for high dimensional data via recursive Lanczos bisection. J. Mach. Learn. Res., 10:1989–2012, 2009. ISSN 1532-4435.

Chan:1985:MAH

[CS85a] T. F. Chan and Y. Saad. Multigrid algorithms on the hypercube multiprocessor. Research Report 368, Department of Computer Science, Yale University, New Haven, CT, USA, February 1985.

Chan:1985:IMS

[CS85b] Tony F. Chan and Youcef Saad. Iterative methods for solving bordered systems with applications to continuation methods. SIAM Journal on Scientific and Statistical Computing, 6(2):438– 451, April 1985. CODEN SI-JCD4. ISSN 0196-5204.

Chan:1986:MAH

[CS86] T. F. Chan and Y. Saad. Multigrid algorithms on the hypercube multiprocessor. IEEE Trans. Comput., C-35(11):969– 977, November 1986. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        1676698.

Cai:1992:ODD

[CS92] X. C. Cai and Y. Saad. Overlapping domain decomposition algorithms for general sparse matrices. Technical report, Army High Performance Computing Research Center, 1992. In preparation.

Cai:1993:ODD

[CS93] X. C. Cai and Y. Saad. Overlapping domain decomposition algorithms for general sparse matrices. Technical Report 93-027, Army High Performance Computing Research Center, Minneapolis, MN, USA, 1993. submitted.

Chow:1994:AIP

[CS94] E. Chow and Y. Saad. Approximate inverse preconditioners for general sparse matrices. Technical Report UMSI 94-101, University of Minnesota Supercomputer Institute, Minneapolis, MN 55415, USA, May 1994.

Cai:1996:ODD

[CS96] Xiao-Chuan Cai and Yousef Saad. Overlapping domain decomposition algorithms for general sparse matrices. Numerical Linear Algebra with Applications, 3(3):221–237, May/June 1996. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (elec-

tronic). URL http://www3.
interscience.wiley.com/cgi-
bin/abstract?ID=15000988.

Castillo:1997:TSA

[CS97a] Paul Castillo and Yousef Saad.

Tensor sum approximation preconditioners. In Proceedings of the Eighth SIAM Conference on Parallel Processing for Scientific Computing (Minneapolis, MN, 1997), page 8. SIAM Press, Philadelphia, PA, USA, 1997.

Chapman:1997:DAK

[CS97b] Andrew Chapman and Yousef Saad. Deflated and augmented Krylov subspace techniques. Numerical Linear Algebra with Applications, 4(1):43– 66, January/February 1997. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (elec-

                                        tronic). URL http://www3.
                                        interscience.wiley.com/cgi-
                                        bin/abstract?ID=15001011;
                                        http://www3.interscience.wiley.
                                        com/cgi-bin/fulltext?ID=15001011&
                                        PLACEBO=IE.pdf.

Chow:1997:IPS

[CS97c] Edmond Chow and Yousef Saad. ILUS: an incomplete LU preconditioner in sparse skyline format. International Journal for Numerical Methods in Fluids, 25(7):739–748, 1997. CO-DEN IJNFDW. ISSN 0271-2091 (print), 1097-0363 (electronic).

Chow:1997:AIT

[CS97d] Edmond Chow and Yousef Saad. Approximate inverse techniques for block-partitioned matrices. SIAM Journal on Scientific Computing, 18(6):1657– 1675, November 1997. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/28157.

Chow:1997:ESI

[CS97e] Edmond Chow and Yousef Saad. Experimental study of ILU preconditioners for indefinite matrices. Journal of Computational and Applied Mathematics, 86(2):387–414, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Chow:1997:PAI

[CS97f] Edmond Chow and Yousef Saad. Parallel approximate inverse preconditioners. In Proceedings of the Eighth SIAM Conference on Parallel Processing for Scientific Computing (Minneapolis, MN, 1997), page 10. SIAM Press, Philadelphia, PA, USA, 1997.

Castillo:1998:PME

[CS98a] Paul Castillo and Yousef Saad.

Preconditioning the matrix exponential operator with applications. Journal of Scientific Computing, 13(3):275–302, 1998. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

Chow:1998:AIP

[CS98b] Edmond Chow and Yousef Saad.

Approximate inverse preconditioners via sparse-sparse iterations. SIAM Journal on Scientific Computing, 19(3):995–1023, May 1998. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/27041.

Calvez:1999:MKA

[CS99] Caroline Le Calvez and Yousef Saad. Modified Krylov acceleration for parallel environments. Applied Numerical Mathematics: Transactions of IMACS, 30(2–3): 191–212, June 10, 1999. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

URL http://www.elsevier.
com/cgi-bin/cas/tree/store/
apnum/cas_sub/browse/browse.
cgi?year=1999&volume=30&issue=

2-3&aid=975. Iterative methods and preconditioners (Berlin, 1997).

Chen:2008:TSO

[CS08] Jie Chen and Yousef Saad. On the tensor SVD and the optimal low rank orthogonal approximation of tensors. SIAM Journal on Matrix Analysis and Applications, 30(4):1709–1734, 2008. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Chen:2009:LVV

[CS09a] Jie Chen and Yousef Saad. Lanczos vectors versus singular vectors for effective dimension reduction. IEEE Transactions on Knowledge and Data Engineering, 21(8):1091–1103, ???? 2009. CODEN ITKEEH. ISSN 1041-4347 (print), 1558-2191 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=4674352.

Chen:2009:TSO

[CS09b] Jie Chen and Yousef Saad. On the tensor SVD and the optimal low rank orthogonal approximation of tensors. SIAM Journal on Matrix Analysis and Applications, 30(4):1709–1734, ???? 2009. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Chen:2012:DSE

[CS12] Jie Chen and Yousef Saad. Dense subgraph extraction with application to community detection. IEEE Transactions on Knowledge and Data Engineering, 24(7):1216–1230, ???? 2012. CODEN ITKEEH. ISSN 1041-4347 (print), 1558-2191 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5677532.

Chow:2014:PKS

[CS14] Edmond Chow and Yousef Saad.

Preconditioned Krylov subspace methods for sampling multivariate Gaussian distributions. SIAM Journal on Scientific Computing, 36(2):A588–A608, ???? 2014. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Chen:2018:PEE

[CS18] Jie Chen and Yousef Saad. A posteriori error estimate for computing tr(f(A)) by using the Lanczos method. Numerical Linear Algebra with Applications, 25(5):??, October 2018. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chan:1985:SEP

[CSS85] T. Chan, Y. Saad, and M. Schultz.

Solving elliptic partial differential equations on the hypercube multiprocessor. Technical Report YALEU/DCS/RR-373, Department of Computer Science, Yale University, New Haven, CT, USA, March 1985.

Chan:1987:SEP

[CSS87] Tony F. Chan, Youcef Saad, and Martin H. Schultz. Solving elliptic partial differential equations on the hypercube multiprocessor. Applied Numerical Mathematics: Transactions of IMACS, 3(??):81–88, 1987. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cai:2002:PIM

[CSS02] X. Cai, Y. Saad, and M. Sosonkina. Parallel iterative methods in modern physical applications. In Computational science—ICCS 2002, Part II (Amsterdam), volume 2330 of Lecture Notes in Computer Science, pages 345–

354. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002.

URL http://link.springer-
ny.com/link/service/series/
0558/bibs/2330/23300345.htm;
 http://link.springer-ny.com/
link/service/series/0558/papers/
2330/23300345.pdf.

Chapman:2000:HOI

[CSW00] Andrew Chapman, Yousef Saad, and Larry Wigton. High-order ILU preconditioners for CFD problems. International Journal for Numerical Methods in Fluids, 33(6):767–788, 2000. CO-DEN IJNFDW. ISSN 0271-2091 (print), 1097-0363 (electronic).

Chen:2022:GCS

[CSZ22] Jie Chen, Yousef Saad, and Zechen Zhang. Graph coarsening: from scientific computing to machine learning. SeMA Journal, 79(1):187–223, 2022. CODEN ???? ISSN 2254-

3902. URL https://doi.org/
10.1007/s40324-021-00282-x.

Chelikowsky:1995:APS

[CTJ^{+}95] James R. Chelikowsky, N. Troullier, X. Jing, D. Dean, N. Binggeli, K. Wu, and Y. Saad. Algorithms for predicting the structural properties of clusters. Computer Physics Communications, 85(3):325–335, March 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/001046559400147T.

Chelikowsky:1993:FDP

[CTS93] James R. Chelikowsky, N. Troullier, and Y. Saad. The finite difference pseudo-potential method: electronic structure calculations without a basis. Technical Report UMSI 93-234, University of Minnesota Supercomputer Institute, Minneapolis, MN 55415, USA, December 1993. appeared in Phys. Rev. Letters, vol. 72 (1994).

Chelikowsky:1994:FDP

[CTS94] James R. Chelikowsky, N. Troullier, and Y. Saad. The finite difference pseudo-potential method: electronic structure calculations without a basis. Physical Review Letters, 72:1240–1243, 1994. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http:

//link.aps.org/doi/10.1103/
PhysRevLett.72.1240.

Chelikowsky:2007:AEE

[CTSZ07] James R. Chelikowsky, Murilo L. Tiago, Yousef Saad, and Yunkai Zhou. Algorithms for the evolution of electronic properties in nanocrystals. Computer Physics Communications, 177 (1–2):1–5, July 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0010465507000458.

Chelikowsky:1994:HOF

[CTWS94] James R. Chelikowsky, N. Troullier, K. Wu, and Y. Saad. Higherorder finite-difference pseudopotential method: An application to diatomic molecules. Physical Review B: Condensed Matter and Materials Physics, 50(16):11355– 11364, ???? 1994. CODEN PRB-MDO. ISSN 1098-0121. URL

                                        http://link.aps.org/doi/10.
                                        1103/PhysRevB.50.11355.

Chelikowsky:2009:AEV

[CZC^{+}09] James R. Chelikowsky, Alexey T. Zayak, T-L Chan, Murilo L. Tiago, Yunkai Zhou, and Yousef Saad. Algorithms for the electronic and vibrational properties of nanocrystals. Journal of Physics: Condensed Matter, 21 (6):064207, February 11, 2009. CODEN JCOMEL. ISSN 0953-8984 (print), 1361-648x (electronic).

Dillon:2018:HLR

[DKXS18] Geoffrey Dillon, Vassilis Kalantzis, Yuanzhe Xi, and Yousef Saad. A hierarchical low rank Schur complement preconditioner for indefinite linear systems. SIAM Journal on Scientific Computing, 40 (4):A2234–A2252, ???? 2018. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

delaGrandmaison:2005:ECC

[dlGGS^{+}05] Emmanuel Lorin de la Grandmaison, Shivaraju B. Gowda, Yousef Saad, Murilo L. Tiago, and James R. Chelikowsky. Efficient computation of the coupling matrix in time-dependent density functional theory. Computer Physics Communications, 167(1):7–22, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

DiNapoli:2016:EEE

[DPS16] Edoardo Di Napoli, Eric Polizzi, and Yousef Saad. Efficient estimation of eigenvalue counts in an interval. Numerical Linear Algebra with Applications, 23(4): 674–692, August 2016. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Datta:1991:SLL

[DS91a] B. N. Datta and Y. Saad. Solution of large linear algebra problems in control. Presentation at the 1986 SIAM meeting on Linear Algebra and its Applications, Boston Mass., Aug 1986, 1991.

Datta:1991:AML

[DS91b] Biswa Nath Datta and Youcef Saad. Arnoldi methods for large Sylvester-like observer problems, and an associated algorithm for partial spectrum assignment. Linear Algebra and its Applications, 154/156:225–244, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

El-Guide:2020:RAM

[EGMS20] Mohamed El-Guide, Agnieszka Miȩdlar, and Yousef Saad. A rational approximation method for solving acoustic nonlinear eigenvalue problems. Engineering Analysis with Boundary Elements, 111:44–54, 2020. ISSN 0955-7997 (print), 1873-197x (electronic).

Elman:1986:HCK

[ESS86] Howard C. Elman, Youcef Saad, and Paul E. Saylor. A hybrid Chebyshev Krylov subspace algorithm for solving nonsymmetric systems of linear equations. SIAM Journal on Scientific and Statistical Computing, 7(3):840– 855, July 1986. CODEN SIJCD4. ISSN 0196-5204.

Fitzgibbon:1986:MCM

[Fit86] W. E. Fitzgibbon, editor. Mathematical and computational methods in seismic exploration and reservoir modeling: [proceedings of the Conference on mathematical and computational methods in seismic exploration and reservoir modeling; Houston, Texas, January 21–24, 1985]. SIAM Press, Philadelphia, PA, USA, 1986. ISBN 0-89871-205-x. LCCN TN269 .C654 1985.

Ferreira:1996:PAI

[FRSY96] A. Ferreira, J. Rolim, Y. Saad, and T. Yang, editors. Parallel algorithms for irregularly structured problems, volume 1117 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-61549-0. LCCN QA76.642 .I59 1996.

Fan:2020:SAP

[FSUS20] Tiffany Fan, David I. Shuman, Shashanka Ubaru, and Yousef Saad. Spectrum-adapted polynomial approximation for matrix functions with applications in graph signal processing. Algorithms (Basel), 13(11):Paper No. 295, 22, November 2020. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https:/

/www.mdpi.com/1999-4893/13/
11/295.

Ferng:1992:SMC

[FWPS92] W. Ferng, K. Wu, S. Petiton, and Y. Saad. Sparse matrix computations on the massively parallel computers. Technical Report 92-084, Army High Performance Computing Research Center, Minneapolis, MN, 1992. (A short version of this report appeared in International Journal of Modern Physics).

Golub:1994:RAI

[GGL94] Gene Golub, Anne Greenbaum, and Mitchell Luskin, editors. Recent advances in iterative methods: [papers from the IMA Workshop on Iterative Methods for Sparse and Structured Problems, held in Minneapolis, Minnesota, February 24–March 1, 1992], volume 60 of The IMA Volumes in Mathematics and its Applications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 0-387-94252-1 (New York), 3-540-94252-1 (Berlin). LCCN QA297.8 .R43 1994.

URL http://zbmath.org/?q=
an:0790.00015.

Giraud:2010:SAS

[GHS10] L. Giraud, A. Haidar, and Y. Saad. Sparse approximations of the Schur complement for parallel algebraic hybrid solvers in 3D. Numerical Mathematics. Theory, Methods and Applications, 3(3):276–294, 2010. CO-DEN ???? ISSN 1004-8979 (print), 2079-7338 (electronic).

Gear:1981:ISL

[GS81] C. W. Gear and Y. Saad. Iterative solution of linear equations in ODE codes. Report UIUCDCS-R-81-1054, Department of Computer Science, University of Illinois at Urbana-Champaign, Urbana, IL, USA, January 1981. 33 pp.

Gear:1983:ISL

[GS83] C. W. Gear and Y. Saad. Iterative solution of linear equations in ODE codes. SIAM Journal on Scientific and Statistical Computing, 4(4):583–601, December 1983. CODEN SIJCD4. ISSN 0196-5204.

Gallopoulos:1987:PBC

[GS87] E. J. (Efstratios J.) Gallopoulos and Y. Saad. A parallel block cyclic reduction algorithm for the fast solution of elliptic equations. Technical Report CSRD 659, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. 14 pp.

Gallopoulos:1988:PBCb

[GS88a] E. Gallopoulos and Y. Saad. A parallel block cyclic reduction algorithm for the fast solution of elliptic equations. In Supercomputing (Athens, 1987), volume 297 of Lecture Notes in Computer Science, pages 563–575. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1988.

Gallopoulos:1988:PBCa

[GS88b] E. J. (Efstratios J.) Gallopoulos and Y. Saad. A parallel block cyclic reduction algorithm for the fast solution of elliptic equations. Technical Report CSRD 753, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 17 + 7 pp.

Gallopoulos:1989:PSPb

[GS89a] E. Gallopoulos and Y. Saad. On the parallel solution of parabolic equations. In R. De Groot, editor, Proceedings of the International Conference on Supercomputing 1989, Heraklion, Crete, June 5–9, 1989. ACM press, 1989.

Gallopoulos:1989:PBC

[GS89b] E. Gallopoulos and Y. Saad. A parallel block cyclic reduction algorithm for the fast solution of elliptic equations. Parallel Computing, 10(2):143–159, April 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gallopoulos:1989:PSPa

[GS89c] E. J. (Efstratios J.) Gallopoulos and Y. Saad. On the parallel solution of parabolic equations. Technical Report CSRD 854, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, June 1989. 22 pp.

Gallopoulos:1989:SFE

[GS89d] E. J. (Efstratios J.) Gallopoulos and Youcef Saad. Some fast elliptic solvers on parallel architectures and their complexities. Technical Report CSRD 862, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, March 1989. 29 pp.

Gallopoulos:1990:PSP

[GS90a] E. Gallopoulos and Y. Saad. Parallel solution of parabolic equations by the rational approximation approach. Technical report, RIACS, NASA Ames research center, Moffett Field, CA, 1990. In preparation.

Gallopoulos:1990:ESP

[GS90b] E. J. (Efstratios J.) Gallopoulos and Y. Saad. Efficient solution of parabolic equations by polynomial approximation methods. Technical Report CSRD 969, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, February 1990. 32 pp.

Gallopoulos:1992:ESPb

[GS92a] E. Gallopoulos and Y. Saad. Efficient solution of parabolic equations by Krylov approximation methods. SIAM Journal on Scientific and Statistical Computing, 13(5):1236–1264, September 1992. CODEN SIJCD4. ISSN 0196-5204.

Gallopoulos:1992:ESPa

[GS92b] E. J. (Efstratios J.) Gallopoulos and Y. Saad. Efficient solution of parabolic equations by Krylov approximation methods. Technical Report CSRD 1147, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 1992. 30 pp.

Goehring:1994:HAA

[GS94] T. Goehring and Y. Saad.

Heuristic algorithms for automatic graph partitioning. Technical Report UMSI 94-29, University of Minnesota Supercomputer Institute, Minneapolis, MN 55415, USA, February 1994. submitted.

Guillaume:2003:RAP

[GSS03] Philippe Guillaume, Yousef Saad, and Masha Sosonkina. Rational approximation preconditioners for sparse linear systems. Journal of Computational and Applied Mathematics, 158(2):419–442, 2003. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Henon:2006:PMI

[HS06] Pascal Hénon and Yousef Saad. A parallel multistage ILU factorization based on a hierarchical graph decomposition. SIAM Journal on Scientific Computing, 28(6):2266–2293, January 2006. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

He:2024:NCN

[HTZ^{+}24] Huan He, Ziyuan Tang, Shifan Zhao, Yousef Saad, and Yuanzhe Xi. nlTGCR: a class of nonlinear acceleration procedures based on conjugate residuals. SIAM Journal on Matrix Analysis and Applications, 45(1):712–743, 2024. CODEN SJMAEL. ISSN 0895-4798,1095-7162. URL https://

                                        doi.org/10.1137/23M1576360.

Ipsen:1985:IPA

[IS85] I. Ipsen and Y. Saad. The impact of parallel architectures on the solution of eigenvalue problems. Technical Report YALEU/DCS/RR-444, Department of Computer Science, Yale University, New Haven, CT, USA, December 1985.

Ipsen:1986:IPAb

[IS86a] I.C.F.Ipsen and Y. Saad. The impact of parallel architectures on large sparse matrix techniques. In R. Willoughby J. Cullum, editor, Proceeding of the 1985 IBM workshop on large eigenvalue computations, Oberlech/Austria, 8-12 July 1985. North-Holland Publishing Co., Amsterdam, The Netherlands, 1986. To appear.

Ipsen:1986:IPAa

[IS86b] I. Ipsen and Y. Saad. The impact of parallel architectures on the solution of eigenvalue problems. In J. Cullum and R. A. Willoughby, editors, Large Scale Eigenvalue Problems. North-Holland, Vol. 127 Mathematics Studies Series, Amsterdam, The Netherlands, 1986.

Ipsen:1984:CDL

[ISS84] I. Ipsen, Y. Saad, and M. H.

Schultz. Complexity of dense linear system solution on a multiprocessor ring. Technical Report 349, Department of Computer Science, Yale University, New Haven, CT, USA, 1984.

Ipsen:1986:CDL

[ISS86] Ilse C. F. Ipsen, Youcef Saad, and Martin H. Schultz. Complexity of dense-linear-system solution on a multiprocessor ring. Linear Algebra and its Applications, 77: 205–239, 1986. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Special volume on parallel computing.

Ilic:2007:LSS

[ITS07] M. Ilić, I. W. Turner, and Y. Saad. Linear system solution by null-space approximation and projection (SNAP). Numerical Linear Algebra with Applications, 14(1):61–82, ???? 2007. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Jay:1999:ESC

[JKSC99] Laurent O. Jay, Hanchul Kim, Yousef Saad, and James R. Chelikowsky. Electronic structure calculations for plane-wave codes without diagonalization. Computer Physics Communications, 118(1):21–30, April 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0010465598001921.

Johnsson:1987:ADM

[JSS87] S. Lennart Johnsson, Youcef Saad, and Martin H. Schultz. Alternating direction methods on multiprocessors. SIAM Journal on Scientific and Statistical Computing, 8(5):686–700, September 1987. CODEN SI-JCD4. ISSN 0196-5204.

Jones:2007:CBI

[JSS07] J. Jones, M. Sosonkina, and Y. Saad. Component-based iterative methods for sparse linear systems. Concurrency and Computation: Practice and Experience, 19(5):625–635, April 10, 2007. CODEN CCPEBO.

ISSN 1532-0626 (print), 1532-0634 (electronic).

Jing:1994:IMD

[JTD^{+}94] Xiaodun Jing, N. Troullier, David Dean, N. Binggeli, James R. Chelikowsky, K. Wu, and Y. Saad. Ab initio moleculardynamics simulations of Si clusters using the higher-order finite-difference-pseudopotential method. Physical Review B: Condensed Matter and Materials Physics, 50(16):12234–12237, ???? 1994. CODEN PRBMDO. ISSN 1098-0121. URL http:

//link.aps.org/doi/10.1103/
PhysRevB.50.12234.

Kokiopoulou:2009:TOE

[KCS09] E. Kokiopoulou, J. Chen, and Y. Saad. Trace optimization and eigenproblems in dimension reduction methods. Technical Report UMSI-2009-31, Minnesota Supercomputer Institute, University of Minnesota, Minneapolis, MN, USA, 2009.

Kokiopoulou:2011:TOE

[KCS11] E. Kokiopoulou, J. Chen, and Y. Saad. Trace optimization and eigenproblems in dimension reduction methods. Numerical Linear Algebra with Applications, 18 (3):565–602, May 2011. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kalantzis:2018:DDA

[KKPS18] Vassilis Kalantzis, James Kestyn, Eric Polizzi, and Yousef Saad. Domain decomposition approaches for accelerating contour integration eigenvalue solvers for symmetric eigenvalue problems. Numerical Linear Algebra with Applications, 25(5):??, October 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kalantzis:2016:SSC

[KLS16] Vassilis Kalantzis, Ruipeng Li, and Yousef Saad. Spectral Schur complement techniques for symmetric eigenvalue problems. Electronic Transactions on Numerical Analysis, 45:305–329, 2016. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic).

URL http://etna.mcs.kent.
edu/vol.45.2016/pp305-329.
dir/pp305-329.pdf;   http:/
/etna.mcs.kent.edu/volumes/
2011-2020/vol45/abstract.php?
vol=45&pages=305-329.

Kalantzis:2018:SID

[KMB^{+}18] Vassilis Kalantzis, A. Cristiano I. Malossi, Costas Bekas, Alessandro Curioni, Efstratios Gallopoulos, and Yousef Saad. A scalable iterative dense linear system solver for multiple right-hand sides in data analytics. Parallel Computing, 74(??):136–153, ???? 2018. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0167819117302119.

Kaagstrom:1983:MPP

[KR83] Bo Kågström and Axel Ruhe, editors. Matrix Pencils: Proceedings of a Conference Held at Pite Havsbad, Sweden, March 22–24, 1982, volume 973 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CO-DEN LNMAA2. ISBN 0-387-11983-3 (print), 3-540-11983-3 (print), 3-540-39447-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 973. URL http:/

/link.springer.com/book/10.
1007/BFb0062089;   http://
www.springerlink.com/content/
978-3-540-39447-1.

Kerkhoven:1987:ATD

[KS87] T. Kerkhoven and Y. Saad. Acceleration techniques for decoupling algorithms in semiconductor simulation. Technical Report 684, University of Illinois, CSRD, Urbana, IL., 1987.

Kerkhoven:1992:AMC

[KS92] T. Kerkhoven and Y. Saad. On acceleration methods for coupled nonlinear elliptic systems. Numerische Mathematik, 60(4):525– 548, January 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kokiopoulou:2005:FRU

[KS05a] Effrosyni Kokiopoulou and Yousef Saad. Face recognition using OPRA-faces. In Proceedings. Fourth International Conference on Machine Learning and Applications, 2005, page ?? IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2005. URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1607433.

Kokiopoulou:2005:ONP

[KS05b] Effrosyni Kokiopoulou and Yousef Saad. Orthogonal neighborhood preserving projections. In Fifth IEEE International Conference on Data Mining, page ?? IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2005. URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1565684.

Kokiopoulou:2007:ONP

[KS07] Effrosyni Kokiopoulou and Yousef Saad. Orthogonal neighborhood preserving projections: a projection-based dimensionality reduction technique. IEEE Transactions on Pattern Analysis and Machine Intelligence, 29 (12):2143–2156, December 2007. CODEN ITPIDJ. ISSN 0162-8828, 1939-3539, 2160-9292.

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
4359299.

Kechroud:2003:PTS

[KSS03] Riyad Kechroud, Azzeddine Soulaimani, and Yousef Saad. Preconditioning techniques for the solution of the Helmholtz equation by the finite element method. In Computational science and its applications— ICCSA 2003. Part II, volume 2668 of Lecture Notes in Computer Science, pages 847–858. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003.

Kechroud:2004:PTS

[KSSG04] Riyad Kechroud, Azzeddine Soulaimani, Yousef Saad, and Shivaraju Gowda. Preconditioning techniques for the solution of the Helmholtz equation by the finite element method. Mathematics and Computers in Simulation, 65(4-5):303–321, 2004. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Kalantzis:2018:BAM

[KXS18] Vassilis Kalantzis, Yuanzhe Xi, and Yousef Saad. Beyond automated multilevel substructuring: Domain decomposition with rational filtering. SIAM Journal on Scientific Computing, 40 (4):C477–C502, ???? 2018. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Little:2002:PPS

[LLCS02] L. Little, Z. Li, H. G. Choi, and Y. Saad. Particle partitioning strategies for the parallel computation of solid-liquid flows. Computers and Mathematics with Applications, 43(12): 1591–1616, June 2002. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Little:2003:BPSa

[LS03] Leigh Little and Yousef Saad. Block preconditioners for saddle point problems. Numerical Algorithms, 33(1–4):367–379, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/
content/getfile/5058/46/30/
abstract.htm;   http://ipsapp007.
kluweronline.com/content/getfile/
5058/46/30/fulltext.pdf. In-

ternational Conference on Numerical Algorithms, Vol. I (Marrakesh, 2001).

Li:2005:CVI

[LS05a] Na Li and Yousef Saad. Crout versions of ILU factorization with pivoting for sparse symmetric matrices. Electronic Transactions on Numerical Analysis, 20:75–85, 2005. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (elec-

tronic). URL http://etna.mcs.
kent.edu/vol.20.2005/pp75-
85.dir/pp75-85.pdf.

Li:2005:SRV

[LS05b] Zhongze Li and Yousef Saad. SchurRAS: A restricted version of the overlapping Schur complement preconditioner. SIAM Journal on Scientific Computing, 27(5):1787–1801, September 2005. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic). URL http:

//epubs.siam.org/volume-27/
art_60835.html.

Li:2006:MMI

[LS06] Na Li and Yousef Saad. MIQR:

A multilevel incomplete QR preconditioner for large sparse leastsquares problems. SIAM Journal on Matrix Analysis and Applications, 28(2):524–550, January 2006. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Li:2013:DCL

[LS13a] Ruipeng Li and Yousef Saad. Divide and conquer low-rank preconditioners for symmetric matrices. SIAM Journal on Scientific Computing, 35(4):A2069– A2095, ???? 2013. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Li:2013:GAP

[LS13b] Ruipeng Li and Yousef Saad.

GPU-accelerated preconditioned iterative linear solvers. The Journal of Supercomputing, 63(2): 443–466, February 2013. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://link.springer.
com/article/10.1007/s11227-
012-0825-3.

Li:2017:LRC

[LS17] Ruipeng Li and Yousef Saad.

Low-rank correction methods for algebraic domain decomposition preconditioners. SIAM Journal on Matrix Analysis and Applications, 38(3):807–828, 2017. CO-DEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Li:2003:CVG

[LSC03] Na Li, Yousef Saad, and Edmond Chow. Crout versions of ILU for general sparse matrices. SIAM Journal on Scientific Computing, 25(2):716–728, March 2003. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/40509.

Lee:1986:EMS

[LSS86] D. Lee, Y. Saad, and M. H. Schultz. An efficient method with for solving the three-dimensional wide angle wave equation. Technical Report YALEU/DCS/RR-463, Department of Computer Science, Yale University, New Haven, CT, USA, 1986.

Li:2003:PPV

[LSS03a] Zhongze Li, Yousef Saad, and Masha Sosonkina. pARMS: a parallel version of the algebraic recursive multilevel solver. Numerical Linear Algebra with Applications, 10(5–6):485–509, July/September 2003. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). Preconditioning, 2001 (Tahoe City, CA).

Little:2003:BPSb

[LSS03b] Leigh Little, Yousef Saad, and Laurent Smoch. Block LU preconditioners for symmetric and nonsymmetric saddle point problems. SIAM Journal on Scientific Computing, 25(2):729–748, March 2003. CODEN SJOCE3.

ISSN 1064-8275 (print), 1095-7197 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/40551.

Lin:2016:ASD

[LSY16] Lin Lin, Yousef Saad, and Chao Yang. Approximating spectral densities of large matrices. SIAM Review, 58(1):34–65, 2016. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Li:2019:ESL

[LXES19] Ruipeng Li, Yuanzhe Xi, Lucas Erlandson, and Yousef Saad. The Eigenvalues Slicing Library (EVSL): Algorithms, implementation, and software. SIAM Journal on Scientific Computing, 41 (4):C393–C415, ???? 2019. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Li:2016:SCB

[LXS16] Ruipeng Li, Yuanzhe Xi, and Yousef Saad. Schur complementbased domain decomposition preconditioners with low-rank corrections. Numerical Linear Algebra with Applications, 23(4): 706–729, August 2016. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Liu:2020:STD

[LXSdH20] Xiao Liu, Yuanzhe Xi, Yousef Saad, and Maarten V. de Hoop. Solving the three-dimensional high-frequency Helmholtz equation using contour integration and polynomial preconditioning.

SIAM Journal on Matrix Analysis and Applications, 41(1):58– 82, 2020. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Li:2016:TRL

[LXV^{+}16] Ruipeng Li, Yuanzhe Xi, Eugene Vecharynski, Chao Yang, and Yousef Saad. A thickrestart Lanczos algorithm with polynomial filtering for Hermitian eigenvalue problems. SIAM Journal on Scientific Computing, 38(4):A2512–A2534, ???? 2016. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

MacLachlan:2012:MCS

[MOKS12] S. MacLachlan, D. Osei-Kuffuor, and Yousef Saad. Modification and compensation strategies for threshold-based incomplete factorizations. SIAM Journal on Scientific Computing, 34 (1):A48–A75, ???? 2012. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

URL http://epubs.siam.org/
sisc/resource/1/sjoce3/v34/
i1/pA48_s1.

Ma:1992:BAP

[MS92] S. Ma and Y. Saad. Block-ADI preconditioners for solving sparse nonsymmetric linear systems of equations. Technical Report 92-161, Minnesota Supercomputer Institute, University of Minnesota, Minneapolis, 1992. Appeared in Numerical Linear Algebra, L. Reichel and A. Ruttan and R. S. Varga eds, Walter de Gruyter publications, New-York,pp. 165-178 (1993).

Ma:1993:BAP

[MS93] Sangback Ma and Youcef Saad.

Block-ADI preconditioners for solving sparse nonsymmetric linear systems of equations. In L. Reichel, A. Ruttan, and R. S. Varga, editors, Numerical linear algebra (Kent, OH, 1992), pages 165–178. Walter de Gruyter, New York, 1993.

Ma:1994:DIS

[MS94] S. Ma and Y. Saad. Distributed ILU(0) and SOR preconditioners for unstructured sparse linear systems. Technical Report 94— , Army High Performance Computing Research Center, University of Minnesota, Minneapolis, MN, 1994.

MacLachlan:2007:GSC

[MS07a] S. MacLachlan and Yousef Saad.

A greedy strategy for coarsegrid selection. SIAM Journal on Scientific Computing, 29(5): 1825–1853, ???? 2007. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

MacLachlan:2007:GCS

[MS07b] Scott MacLachlan and Yousef Saad. Greedy coarsening strategies for nonsymmetric problems. SIAM Journal on Scientific Computing, 29(5):2115– 2143, ???? 2007. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Ngo:2010:TRO

[NBS10] T. T. Ngo, M. Bellalij, and Y. Saad. The trace ratio optimization problem for dimensionality reduction. SIAM Journal on Matrix Analysis and Applications, 31(5):2950– 2971, ???? 2010. CO-DEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/simax/resource/1/
                                        sjmael/v31/i5/p2950_s1.

Ngo:2012:TRO

[NBS12] T. T. Ngo, M. Bellalij, and Y. Saad. The trace ratio optimization problem. SIAM Review, 54(3):545–569, ???? 2012. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Napoli:2016:EEE

[NPS16] Edoardo Di Napoli, Eric Polizzi, and Yousef Saad. Efficient estimation of eigenvalue counts in an interval. Numerical Linear Algebra with Applications, 23(4):674– 692, 2016. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ogut:2003:ICL

[ÖBSC03] Serdar Öğüt, Russ Burdick, Yousef Saad, and James R. Chelikowsky. Ab initio calculations for large dielectric matrices of confined systems. Physical Review Letters, 90(12):127401:1– 127401:??, March 28, 2003. CO-DEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http:

//link.aps.org/doi/10.1103/
PhysRevLett.90.127401.

Osei-Kuffuor:2015:MRU

[OKLS15] Daniel Osei-Kuffuor, Ruipeng Li, and Yousef Saad. Matrix reordering using multilevel graph coarsening for ILU preconditioning. SIAM Journal on Scientific Computing, 37(1):A391– A419, 2015. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Osei-Kuffuor:2010:PHL

[OKS10] Daniel Osei-Kuffuor and Yousef Saad. Preconditioning Helmholtz linear systems. Applied Numerical Mathematics: Transactions of IMACS, 60(4):420–431, April 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Parlett:1985:CSI

[PS85] Beresford N. Parlett and Y. Saad.

Complex shift and invert strategies for real matrices. Research report RR-424, Department of Computer Science, Yale University, New Haven, CT, USA, 1985. 17 pp.

Parlett:1987:CSI

[PS87] Beresford N. Parlett and Youcef Saad. Complex shift and invert strategies for real matrices. Linear Algebra and its Applications, 88/89(??):575– 595, 1987. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        https://www.math.utah.edu/
                                        pub/bibnet/authors/s/saad-
                                        yousef.bib;   https://www.
                                        math.utah.edu/pub/tex/bib/
                                        linala1980.bib.

Philippe:1989:SLS

[PS89] B. Philippe and Y. Saad. Solving large sparse eigenvalue problems on supercomputers. In Proceedings of International Workshop on Parallel Algorithms and Architectures, Bonas, France Oct. 3-6 1988. North-Holland Publishing Co., Amsterdam, The Netherlands, 1989.

Philippe:2007:CED

[PS07] Bernard Philippe and Yousef Saad. On correction equations and domain decomposition for computing invariant subspaces. Computer Methods in Applied Mechanics and Engineering, 196 (8):1471–1483, 2007. CODEN CMMECC. ISSN 0045-7825, 0374-2830.

Polizzi:2020:CMS

[PS20] Eric Polizzi and Yousef Saad. Computational materials science and engineering. In Parallel Algorithms in Computational Science and Engineering, Model. Simul. Sci. Eng. Technol., pages 123–150. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2020.

Philippe:1992:NMM

[PSS92] B. Philippe, Y. Saad, and W. J. Stewart. Numerical methods in Markov chain modeling. Journal of Operations Research, 40: ??, 1992.

Petiton:1993:BSM

[PSWF93] S. Petiton, Y. Saad, K. Wu, and W. Ferng. Basic sparse matrix computations on the CM-5. Internat. J. of Modern Physics, 4: 65–83, 1993.

Fang:2009:TCM

[rFS09] Haw ren Fang and Yousef Saad. Two classes of multisecant methods for nonlinear acceleration. Numerical Linear Algebra with Applications, 16(3):197–221, ???? 2009. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Fang:2012:FLP

[rFS12] Haw ren Fang and Yousef Saad.

A filtered Lanczos procedure for extreme and interior eigenvalue problems. SIAM Journal on Scientific Computing, 34(4):A2220– A2246, ???? 2012. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Rocca:2008:TCT

[RGSB08] Dario Rocca, Ralph Gebauer, Yousef Saad, and Stefano Baroni. Turbo charging timedependent density-functional theory with Lanczos chains. Journal of Chemical Physics, 128 (15):154105, April 21, 2008. CO-DEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic).

Reichel:1993:NLA

[RRV93] Lothar Reichel, Arden Ruttan, and Richard S. Varga, editors. Numerical linear algebra: proceedings of the Conference in Numerical Linear Algebra and Scientific Computation, Kent (Ohio), USA, March 13– 14, 1992. Walter de Gruyter, New York, 1993. ISBN 3-11-013784-4. LCCN QA184.C655 1993.

Ruhe:1994:RKA

[Ruh94]  Axel Ruhe.  The rational

Krylov algorithm for nonsymmetric eigenvalue problems. III: Complex shifts for real matrices. BIT (Nordisk tidskrift for informationsbehandling), 34(1):165– 176, March 1994. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL http://www.mai.liu.se/
BIT/contents/bit34.html;
http://www.springerlink.com/
openurl.asp?genre=article&
issn=0006-3835&volume=34&issue=
1&spage=165.

Saad:1974:CEL

[Saa74a] Y. Saad. Computation of eigenvalues of large Hermitian matrices by partitioning techniques. Doctorate de 3ème cycle, INPG - University of Grenoble, Grenoble, France, 1974.

Saad:1974:ETO

[Saa74b] Youcef Saad. Études des translations d’origine dans les algorithmes LR et QR. C. R. Acad.

Sci. Paris Sér. A, 278:93–96, 1974.

Saad:1974:SOA

[Saa74c] Youcef Saad. Shifts of origin for the OR algorithm. In Information processing 74 (Proc. IFIP Congress, Stockholm, 1974), pages 527–531. North-Holland Publishing Co., Amsterdam, The Netherlands, 1974.

Saad:1980:LBA

[Saa80a] Y. Saad. The Lanczos biorthogonalization algorithm and other oblique projection methods for solving large unsymmetric systems. Report 1036, Department of Computer Science, University of Illinois at Urbana-Champaign, Urbana, IL, USA, 1980. 44 pp.

Saad:1980:RCL

[Saa80b] Y. Saad. On the rates of convergence of the Lanczos and the block Lanczos methods. SIAM Journal on Numerical Analysis, 17(5):687–706, October 1980. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Saad:1980:VAM

[Saa80c] Y. Saad. Variations on Arnoldi’s method for computing eigenelements of large unsymmetric matrices. Linear Algebra and its Applications, 34:269–295, 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Saad:1981:KSM

[Saa81] Y. Saad. Krylov subspace methods for solving large unsymmetric linear systems. Mathematics of Computation, 37(155):105– 126, 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Saad:1982:LBA

[Saa82a] Y. Saad. The Lanczos biorthogonalization algorithm and other oblique projection methods for solving large unsymmetric systems. SIAM Journal on Numerical Analysis, 19(3):485–506, June 1982. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Saad:1982:PMS

[Saa82b] Y. Saad. Projection methods for solving large sparse eigenvalue problems. In B. Kagstrom and A. Ruhe, editors, Matrix Pencils, proceedings, Pitea Havsbad, pages 121–144. University of Umea, Sweden, Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. Lecture notes in Math. Series, Number 973.

Saad:1983:LSP

[Saa83a] Y. Saad. Least squares polynomials in the complex plane with applications to solving sparse nonsymmetric matrix problems. Technical report YALEU/DCS/RR-276, Department of Computer Science, Yale University, New Haven, CT, USA, 1983.

Saad:1983:NMS

[Saa83b] Y. Saad. Numerical methods for the solution of large dimension matrix problems. Doctorat d’etat, Université scientifique et medicale de Grenoble et INPG, Grenoble, France, 1983.

Saad:1983:BRB

[Saa83c] Youcef Saad. Book review: The Symmetric Eigenvalue Problem (Beresford N. Parlett). SIAM Review, 25(2):286–287, ???? 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Saad:1983:ISI

[Saa83d] Youcef Saad. Iterative solution of indefinite symmetric linear systems by methods using orthogonal polynomials over two disjoint intervals. SIAM Journal on Numerical Analysis, 20(4):784– 811, August 1983. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Saad:1983:PMS

[Saa83e] Youcef Saad. Projection methods for solving large sparse eigenvalue problems. Lecture Notes in Mathematics, 973:121–144, 1983. CODEN LNMAA2. ISBN 0-387-11983-3, 3-540-11983-3 (print), 3-540-39447-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/
10.1007/BFb0062098/.

Saad:1984:CNM

[Saa84a]  Y. Saad.  On the condi-

tion numbers of modified moment matrices arising in least squares approximation in the complex plane. Technical Report YALEU/DCS/RR-347, Department of Computer Science, Yale University, New Haven, CT, USA, 1984.

Saad:1984:CAT

[Saa84b] Youcef Saad. Chebyshev acceleration techniques for solving nonsymmetric eigenvalue problems. Mathematics of Computation, 42 (166):567–588, April 1984. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Saad:1984:PUS

[Saa84c] Youcef Saad. Practical use of some Krylov subspace methods for solving indefinite and nonsymmetric linear systems. SIAM Journal on Scientific and Statistical Computing, 5(1):203–228, March 1984. CODEN SIJCD4. ISSN 0196-5204.

Saad:1985:CCG

[Saa85a] Y. Saad. Communication complexity of the Gaussian elimination algorithm on multiprocessors. Research Report 348, Department of Computer Science, Yale University, New Haven, CT, USA, 1985.

Saad:1985:PEL

[Saa85b] Y. Saad. Partial eigensolutions of large nonsymmetric matrices. Technical Report YALEU/DCS/RR-397, Department of Computer Science, Yale University, New Haven, CT, USA, 1985.

Saad:1985:PUP

[Saa85c] Youcef Saad. Practical use of polynomial preconditionings for the conjugate gradient method. SIAM Journal on Scientific and Statistical Computing, 6(4):865– 881, October 1985. CODEN SI-JCD4. ISSN 0196-5204.

Saad:1986:GEHa

[Saa86a] Y. Saad. Gaussian elimination on hypercubes. Technical Report YALEU/DCS/RR-462, Department of Computer Science, Yale University, New Haven, CT, USA, 1986.

Saad:1986:CNM

[Saa86b] Y. Saad. On the condition numbers of modified moment matrices arising in least squares approximation in the complex plane. Numerische Mathematik, 48:337–347, 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Saad:1986:CCG

[Saa86c] Youcef Saad. Communication complexity of the Gaussian elimination algorithm on multiprocessors. Linear Algebra and its Applications, 77:315–340, 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Special volume on parallel computing.

Saad:1986:GEHb

[Saa86d] Youcef Saad. Gaussian elimination on hypercubes. In M. Cosnard, Y. Robert, P. Quinton, and M. Tchuente, editors, Parallel algorithms & architectures (Luminy, 1986), pages 5–17. North-Holland Publishing Co., Amsterdam, The Netherlands, 1986.

Saad:1986:CNS

[Saa86e] Youcef Saad. On the condition number of some Gram matrices arising from least squares approximation in the complex plane. Numerische Mathematik, 48(3):337–347, March 1986. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Saad:1987:DPNa

[Saa87a] Y. Saad. On the design of parallel numerical methods in message passing and shared memory environments. Technical Report CSRD-614, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. 20 pp.

Saad:1987:DPNb

[Saa87b] Y. Saad. On the design of parallel numerical methods in message passing and shared memory environments. In ????, editor, Proceedings of International Seminar on Scientific Supercomputers, Paris, France, February 2–6, 1987. ????, ????, 1987.

Saad:1987:LSP

[Saa87c] Youcef Saad. Least squares polynomials in the complex plane and their use for solving nonsymmetric linear systems. SIAM Journal on Numerical Analysis, 24 (1):155–169, February 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Saad:1987:LMS

[Saa87d] Youcef Saad. On the Lanczos method for solving symmetric systems with several righthand sides. Mathematics of Computation, 48(178):651–662, April 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Saad:1988:PTI

[Saa88a] Y. Saad. Preconditioning techniques for indefinite and nonsymmetric linear systems. Journal of Comp. and Appl. Math., 24:89– 105, 1988.

Saad:1988:PTNa

[Saa88b] Y. Saad. Preconditioning techniques for nonsymmetric and indefinite linear systems. Technical Report CSRD 792, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 18 + 3 pp.

Saad:1988:PTNb

[Saa88c] Youcef Saad. Preconditioning techniques for nonsymmetric and indefinite linear systems. Journal of Computational and Applied Mathematics, 24(1–2):89– 105, 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). Iterative methods for the solution of linear systems.

Saad:1988:PDM

[Saa88d] Youcef Saad. Projection and deflation methods for partial pole assignment in linear state feedback. IEEE Transactions on Automatic Control, 33(3):290– 297, 1988. CODEN IETAA9. ISSN 0018-9286 (print), 1558-2523 (electronic).

Saad:1989:KSM

[Saa89a]  Youcef Saad.  Krylov sub-

space methods on supercomputers. SIAM Journal on Scientific and Statistical Computing, 10(6): 1200–1232, November 1989. CO-DEN SIJCD4. ISSN 0196-5204. Sparse matrix algorithms on supercomputers.

Saad:1989:NSL

[Saa89b] Youcef Saad. Numerical solution of large nonsymmetric eigenvalue problems. Computer Physics Communications, 53(1–3):71–90, May 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0010465589901495.

Practical iterative methods for large scale computations (Minneapolis, MN, 1988).

Saad:1990:SBT

[Saa90a] Y. Saad. SPARSKIT: A basic tool kit for sparse matrix computations. Technical Report 90-20, Research Institute for Advanced Computer Science, NASA Ames Research Center, Moffett Field, CA, 1990.

Saad:1990:KSM

[Saa90b] Youcef Saad. Krylov subspace methods: theory, algorithms, and applications. In Computing methods in applied sciences and engineering (Paris, 1990), pages 24–41. SIAM Press, Philadelphia, PA, USA, 1990.

Saad:1990:NSL

[Saa90c] Youcef Saad. Numerical solution of large Lyapunov equations. In M. A. Kaashoek, J. H. van Schuppen, and A. C. Ran, editors, Signal Processing, Scattering, Operator Theory, and Numerical Methods. Proceedings of the International Symposium MTNS-89, vol III (Amsterdam, 1989), volume 5 of Progr. Systems Control Theory, pages 503–511. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1990.

Saad:1990:OKS

[Saa90d] Youcef Saad. An overview of Krylov subspace methods with applications to control problems. In M. A. Kaashoek, J. H. van Schuppen, and A. C. Ran, editors, Signal Processing, Scattering, Operator Theory, and Numerical Methods. Proceedings of the international symposium MTNS-89, vol III, (Amsterdam, 1989), volume 5 of Progr. Systems Control Theory, pages 401–410. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1990.

Saad:1991:FIO

[Saa91a] Y. Saad. A flexible innerouter preconditioned GMRES algorithm. Technical Report 91-279, Minnesota Supercomputer Institute, University of Minnesota, Minneapolis, Minnesota, 1991. Appeared in SISSC, vol.4, 1993.

Saad:1991:SIP

[Saa91b] Y. Saad. Supercomputer implementations of preconditioned Krylov subspace methods. Technical Report 91-311, Minnesota Supercomputer Institute, University of Minnesota, Minneapolis, Minnesota, 1991.

Saad:1991:PMN

[Saa91c] Youcef Saad. Projection methods for the numerical solution of Markov chain models. In Numerical solution of Markov chains, volume 8 of Probab. Pure Appl., pages 455–471. Dekker, New York, 1991.

Saad:1992:AST

[Saa92a] Y. Saad. Algorithms and software tools for general purpose automatic domain decomposition. Technical report, Army High Performance Computing Research Center, 1992. In preparation.

Saad:1992:ASK

[Saa92b] Y. Saad. Analysis of some Krylov subspace approximations to the matrix exponential operator. SIAM Journal on Numerical Analysis, 29(1):209–228, February 1992. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Saad:1992:IPM

[Saa92c] Y. Saad. ILUM: A parallel multielimination ILU preconditioner for general sparse matrices. Technical Report 92-241, University of Minnesota, Army High Performance Computing Research Center, Minneapolis, Minnesota, 1992. submitted, under revision.

Saad:1992:IDT

[Saa92d] Y. Saad. ILUT: A dual threshold incomplete ILU factorization. Technical Report 92-38, Minnesota Supercomputer Institute, University of Minnesota, Minneapolis, 1992. to appear.

Saad:1992:KSM

[Saa92e] Y. Saad. Krylov subspace methods in distributed computing environments. Technical Report 92-126, Army High Performance Computing Research Center, Minneapolis, MN, 1992.

Saad:1992:NTP

[Saa92f] Youcef Saad. New trends in preconditioned Krylov subspace methods. In Proceedings of International Conference on Scientific Computation (Hangzhou, 1991), volume 1 of Ser. Appl. Math., pages 121–136. World Sci. Publ., River Edge, NJ, 1992.

Saad:1992:NML

[Saa92g] Youcef Saad. Numerical methods for large eigenvalue problems. Algorithms and Architectures for Advanced Scientific Computing. Halsted Press, New York, USA, 1992. ISBN 0-7190-3386-1 (UK), 0-470-21820-7 (US). xii + 346 pp. LCCN QA188 .S18 1992.

Saad:1992:RAP

[Saa92h] Youcef Saad. Review of ABS projection algorithms by J. Abaffy and E. Spedicato. Linear Algebra and its Applications, 165:267–??, March 1, 1992. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Saad:1993:FIO

[Saa93a] Youcef Saad. A flexible innerouter preconditioned GMRES algorithm. SIAM Journal on Scientific Computing, 14(2): 461–469, March 1993. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Saad:1993:SIP

[Saa93b] Youcef Saad. Supercomputer implementations of preconditioned Krylov subspace methods. In Algorithmic trends in computational fluid dynamics (1991), ICASE/NASA LaRC Ser., pages 107–136. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993.

Saad:1994:DSA

[Saa94a] Y. Saad. Data structures and algorithms for domain decomposition and distributed sparse matrix computations. Technical Report 94—, Army High Performance Computing Research Center, Minneapolis, MN, USA, 1994.

Saad:1994:TEBa

[Saa94b]  Y. Saad.  Theoretical error

bounds and general analysis of a few Lanczos-type algorithms. Technical Report UMSI 94-90, University of Minnesota Supercomputer Institute, Minneapolis, MN 55415, USA, 1994.

Saad:1994:HPP

[Saa94c] Youcef Saad. Highly parallel preconditioners for general sparse matrices. In G. Golub, M. Luskin, and A. Greenbaum, editors, Recent Advances in Iterative Methods, volume 60 of IMA volumes in Mathematics and its Applications, pages 165–199. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994.

Saad:1994:IDT

[Saa94d] Yousef Saad. ILUT: a dual threshold incomplete LU factorization. Numerical Linear Algebra with Applications, 1(4):387– 402, 1994. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Saad:1994:TEBb

[Saa94e] Yousef Saad. Theoretical error bounds and general analysis of a few Lanczos-type algorithms. In Brown et al. [BCEP94], pages 123–134. ISBN 0-89871-339-0. LCCN QC19.2 .C67 1993.

Saad:1995:BNR

[Saa95] Youcef Saad. Book news & reviews: Concurrent Scientific Computing by Eric F. Van de Velde. IEEE Computational Science & Engineering, 2(1):94, Spring 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Saad:1996:IME

[Saa96]  Y. Saad.  ILUM: A multi-

elimination ILU preconditioner for general sparse matrices. SIAM Journal on Scientific Computing, 17(4):830–847, July 1996. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Saad:1997:AAK

[Saa97] Yousef Saad. Analysis of augmented Krylov subspace methods. SIAM Journal on Matrix Analysis and Applications, 18(2):435–449, April 1997. CO-DEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/29428.

Saad:1998:PKS

[Saa98] Y. Saad. Preconditioned Krylov subspace methods. In Algorithms for large scale linear algebraic systems (Gran Canaria, 1996), volume 508 of NATO Adv. Sci. Inst. Ser. C Math. Phys. Sci., pages 131–149. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1998.

Saad:2000:E

[Saa00a] Yousef Saad. Editorial. Numerical Linear Algebra with Applications, 7(7–8):489–490, October/December 2000. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (elec-

tronic). URL http://www3.
interscience.wiley.com/cgi-
bin/abstract/73505473/START;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=73505473&PLACEBO=IE.pdf.

Saad:2000:FAM

[Saa00b] Yousef Saad. Further analysis of minimum residual iterations. Numerical Linear Algebra with Applications, 7(2):67–93, March 2000. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http://www3.

interscience.wiley.com/cgi-
bin/abstract/71008526/START;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=71008526&PLACEBO=IE.pdf.

Saad:2001:PIM

[Saa01] Y. Saad. Parallel iterative methods for sparse linear systems. In Inherently parallel algorithms in feasibility and optimization and their applications (Haifa, 2000), volume 8 of Stud. Comput. Math., pages 423–440. North-Holland Publishing Co., Amsterdam, The Netherlands, 2001.

Saad:2003:FEA

[Saa03a] Yousef Saad. Finding exact and approximate block structures for ILU preconditioning. SIAM Journal on Scientific Computing, 24(4):1107–1123, July 2003. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/39339.

Saad:2003:IMS

[Saa03b] Yousef Saad. Iterative methods for sparse linear systems. SIAM Press, Philadelphia, PA, USA, second edition, 2003. ISBN 0-89871-534-2. xviii + 528 pp. LCCN QA188 .S17 2003.

Saad:2005:MIR

[Saa05] Yousef Saad. Multilevel ILU with reorderings for diagonal dominance. SIAM Journal on Scientific Computing, 27(3):1032–1057, May 2005. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic). URL http://

                                        epubs.siam.org/SISC/volume-
                                        27/art_60273.html.

Saad:2006:FCR

[Saa06]  Yousef Saad.  Filtered con-

jugate residual-type algorithms with applications. SIAM Journal on Matrix Analysis and Applications, 28(3):845–870, January 2006. CODEN SJMAEL.

ISSN 0895-4798 (print), 1095-7162 (electronic).

Saad:2007:SCP

[Saa07] Yousef Saad. Schur complement preconditioners for distributed general sparse linear systems. In Domain decomposition methods in science and engineering XVI, volume 55 of Lect. Notes Comput. Sci. Eng., pages 127–138. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007.

Saad:2011:KSM

[Saa11a] Youcef Saad. Krylov Subspace Methods, chapter 6, pages 125– 162. Volume 66 of Classics in applied mathematics [Saa11b], second edition, 2011. ISBN 1-61197-072-5. LCCN QA188 .S18 2011.

URL http://www.cs.umn.edu/
~saad/eig_book_2ndEd.pdf.

Saad:2011:NML

[Saa11b] Youcef Saad. Numerical Methods for Large Eigenvalue Problems, volume 66 of Classics in applied mathematics. SIAM Press, Philadelphia, PA, USA, second edition, 2011. ISBN 1-61197-072-5. xv + 276 pp. LCCN QA188 .S18 2011. URL http:

//www.cs.umn.edu/~saad/eig_
book_2ndEd.pdf.

Saad:2016:ASI

[Saa16] Yousef Saad. Analysis of subspace iteration for eigenvalue problems with evolving matrices. SIAM Journal on Matrix Analysis and Applications, 37(1):103– 122, ???? 2016. CODEN SJ-MAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Saad:2020:IML

[Saa20] Yousef Saad. Iterative methods for linear systems of equations: a brief historical journey. In 75 years of mathematics of computation, volume 754 of Contemp. Math., pages 197–215. American Mathematical Society, Providence, RI, USA, 2020.

Saad:2022:ODK

[Saa22] Yousef Saad. The origin and development of Krylov subspace methods. Computing in Science and Engineering, 24(4):28– 39, July/August 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Saad:2023:RBJ

[Saa23] Yousef Saad. Revisiting the (block) Jacobi subspace rotation method for the symmetric eigenvalue problem. Numerical Algorithms, 92(1):917–944, January 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https:/

                                        /link.springer.com/article/
                                        10.1007/s11075-022-01377-w.

Saad:2024:FKC

[Saa24] Yousef Saad. Five key concepts that shaped iterative solution methods for linear systems. SIAM News, 57(2):3, 7, 2024. ISSN 1557-9573.

Saad:2025:AMF

[Saa25]  Yousef Saad.  Acceleration

methods for fixed-point iterations. Acta Numerica, 34: 805–890, July 2025. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

URL https://www.cambridge.
org/core/journals/acta-numerica/
article/acceleration-methods-
for-fixedpoint-iterations/
2FABE7BD50568960C8CD86A6F4D77D72.

Saad:2000:PTL

[SAD^{+}00] Yousef Saad, Owe Axelsson, Iain Duff, Wei-Pai Tang, and Andy Wathen, editors. Preconditioning techniques for large sparse matrix problems in industrial applications. Wiley, New York, NY, USA, 2000. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). Papers from the International Conference (SPARSE ’99) held at the University of Minnesota, Minneapolis, MN, June 10–12, 1999, Numer. Linear Algebra Appl. 7 (2000), no. 7–8.

Saad:2010:NME

[SCS10] Yousef Saad, James R. Chelikowsky, and Suzanne M. Shontz. Numerical methods for electronic structure calculations of materials. SIAM Review, 52(1):3–54, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schofield:2012:SSM

[SCS12] Grady Schofield, James R. Chelikowsky, and Yousef Saad. A spectrum slicing method for the Kohn–Sham problem. Computer Physics Communications, 183 (3):497–505, March 2012. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465511003675.

Salas:2015:SRS

[SGSM15] Pablo Salas, Luc Giraud, Yousef Saad, and Stéphane Moreau. Spectral recycling strategies for the solution of nonlinear eigenproblems in thermoacoustics. Numerical Linear Algebra with Applications, 22(6):1039–1058, 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Said:1988:HES

[SKBS88] M. Said, M. A. Kanesha, M. Balkanski, and Y. Saad. Higher excited states of acceptors in cubic semiconductors. Physical Review B: Condensed Matter and Materials Physics, 35(2): 687–695, 1988. CODEN PRB-MDO. ISSN 1098-0121. URL

http://link.aps.org/doi/10.
1103/PhysRevB.35.687.

Saad:1997:PPL

[SKL^{+}97] Yousef Saad, Sergey Kuznetsov, Gen-Ching Lo, Andrei Malevsky, and Andrew Chapman. P SPARSLIB: a portable library of parallel sparse iterative solvers. In Proceedings of the Eighth SIAM Conference on Parallel Processing for Scientific Computing (Minneapolis, MN, 1997), page 8. SIAM Press, Philadelphia, PA, USA, 1997.

Saad:1986:NAS

[SL86] Y. Saad and D. Lee. A new algorithm for solving the wide angle wave equation. Technical Report YALEU/DCS/RR-Draft, Department of Computer Science, Yale University, New Haven, CT, USA, 1986.

Saad:1988:NAS

[SL88] Youcef Saad and Ding Lee. A new algorithm for solving the wide angle wave equation. In Computational acoustics, Vol. 2 (New Haven, Conn., 1986), pages 119–132. North-Holland Publishing Co., Amsterdam, The Netherlands, 1988.

Shi:2022:NPA

[SLX^{+}22] Jia Shi, Ruipeng Li, Yuanzhe Xi, Yousef Saad, and Maarten V. de Hoop. A non-perturbative approach to computing seismic normal modes in rotating planets. Journal of Scientific Computing, 91(2):Paper No. 67, 52, 2022. CODEN JSCOEB. ISSN 0885-7474,1573-7691. URL https://

doi.org/10.1007/s10915-022-
01836-5.

Saad:1995:DSC

[SM95] Y. Saad and A. V. Malevsky.

Data structures, computational, and communication kernels for distributed memory sparse iterative solvers. Lecture Notes in Computer Science, 964:252– ??, 1995. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic).

Sosonkina:2000:PSL

[SMSW00] Maria Sosonkina, John T. Melson, Yousef Saad, and Layne T. Watson. Preconditioning strategies for linear systems arising in tire design. Numerical Linear Algebra with Applications, 7(7– 8):743–757, October/December 2000. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http://www3.

interscience.wiley.com/cgi-
bin/abstract/73505476/START;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=73505476&PLACEBO=IE.pdf.

Preconditioning techniques for large sparse matrix problems in industrial applications (Minneapolis, MN, 1999).

Stathopoulos:2000:PMT

[SÖS^{+}00] Andreas Stathopoulos, Serdar Ögüt, Yousef Saad, James Chelikowsky, and Hanchul Kim. Parallel methods and tools for predicting material properties. Computing in Science and Engineering, 2(4):19–33, July/August 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

dlib.computer.org/cs/books/
cs2000/pdf/c4019.pdf;   http:
//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=852388;
 http://www.computer.org/cse/
cs1999/c4019abs.htm.

Sakellaridi:2008:GBM

[SrFS08] Sophia Sakellaridi, Haw ren Fang, and Yousef Saad. Graphbased multilevel dimensionality reduction with applications to eigenfaces and latent semantic indexing. In ICMLA ’08. Seventh International Conference on Machine Learning and Applications, 2008, pages 194–200. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2008. URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4724975.

Saad:1980:PBS

[SS80] Y. Saad and A. H. Sameh. A parallel block Stiefel method for solving positive definite systems. In M. H. Schultz, editor, Proc. Elliptic Problem Solver Conf., pages 405–412. Academic Press, New York, NY, USA, 1980.

Saad:1981:IMS

[SS81] Y. Saad and A. Sameh. Iterative methods for the solution of elliptic differential equations on multiprocessors. In Wolfgang Handler, editor, Proceedings of the CONPAR 81 Conference, pages 395–411. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981.

Saad85e

[SS85a] Y. Saad and M. Schultz.

Data communication in hypercubes. Technical Report YALEU/DCS/RR-428, Department of Computer Science, Yale University, New Haven, CT, USA, October 1985.

Saad:1985:TPH

[SS85b] Y. Saad and M. Schultz. Topological properties of hypercubes. Technical Report YALEU/DCS/RR-389, Department of Computer Science, Yale University, New Haven, CT, USA, June 1985.

Saad:1985:ADM

[SS85c] Y. Saad and M. H. Schultz. Alternating direction methods on multiprocessors: An extended abstract. Research Report YALEU/DCS/RR-381, Department of Computer Science, Yale University, New Haven, CT, USA, April 1985.

Saad:1985:DCH

[SS85d] Y. Saad and M. H. Schultz. Data communication in hypercubes. Technical Report 428, Department of Computer Science, Yale University, New Haven, CT, USA, 1985. to appear.

Saad:1985:DPM

[SS85e] Y. Saad and M. H. Schultz. Direct parallel methods for solving banded linear systems. Research Report 387, Department of Computer Science, Yale University, New Haven, CT, USA, 1985.

Saad:1985:PIP

[SS85f] Y. Saad and M. H. Schultz. Parallel implementations of preconditioned conjugate gradient methods. Research Report 425, Department of Computer Science, Yale University, New Haven, CT, USA, 1985.

Saad:1985:CGL

[SS85g] Youcef Saad and Martin H.

Schultz. Conjugate gradient-like algorithms for solving nonsymmetric linear systems. Mathematics of Computation, 44(170):417– 424, April 1985. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Saad:1986:PIP

[SS86a] Y. Saad and M. Schultz. Parallel implementations of preconditioned conjugate gradient methods. In Fitzgibbon [Fit86], page ?? ISBN 0-89871-205-x. LCCN TN269 .C654 1985.

Saad:1986:DCP

[SS86b] Y. Saad and M. H. Schultz.

Data communication in parallel architectures. Research Report YALEU/DCS/RR-461, Department of Computer Science, Yale University, New Haven, CT, USA, 1986.

Saad:1986:GGM

[SS86c] Youcef Saad and Martin H.

Schultz. GMRES: a generalized minimal residual algorithm for solving nonsymmetric linear systems. SIAM Journal on Scientific and Statistical Computing, 7 (3):856–869, July 1986. CODEN SIJCD4. ISSN 0196-5204.

Saad:1987:PDM

[SS87] Y. Saad and M. Schultz. Parallel direct methods for solving banded linear systems. Linear Algebra and its Applications, 88/89(??):623–650, 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Saad:1988:TPH

[SS88] Y. Saad and M. H. Schultz. Topological properties of hypercubes. IEEE Trans. Comput., 37 (7):867–872, 1988. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Saad:1989:DCH

[SS89a] Y. Saad and M. H. Schultz. Data communication in hypercubes. Journal of Parallel and Distributed Computing, 6(1): 115–135, February 1989. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Saad:1989:DCP

[SS89b] Youcef Saad and Martin H. Schultz. Data communication in parallel architectures. Parallel Computing, 11(2):131–150, ???? 4, 1989. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Saad:1998:SDS

[SS98a] Y. Saad and M. Sosonkina. Solution of distributed sparse linear systems using PSPARSLIB. Lecture Notes in Computer Science, 1541:503–509, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stathopoulos:1998:RTJ

[SS98b] Andreas Stathopoulos and Yousef Saad. Restarting techniques for the (Jacobi–)Davidson symmetric eigenvalue methods. Electronic Transactions on Numerical Analysis, 7:163–181, 1998. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (elec-

tronic). URL http://etna.mcs.
kent.edu/vol.7.1998/pp163-
181.dir/pp163-181.pdf. Large

scale eigenvalue problems (Argonne, IL, 1997).

Saad:1999:DSC

[SS99a] Yousef Saad and Maria Sosonkina. Distributed Schur complement techniques for general sparse linear systems. SIAM Journal on Scientific Computing, 21(4):1337–1356, July 1999. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/32899.

Saad:1999:EPM

[SS99b] Yousef Saad and Maria Sosonkina. Enhanced parallel multicolor preconditioning techniques for linear systems. In Proceedings of the Ninth SIAM Conference on Parallel Processing for Scientific Computing 1999 (San Antonio, TX), page 10. SIAM Press, Philadelphia, PA, USA, 1999.

Saad:1999:NSP

[SS99c] Yousef Saad and Maria Sosonkina. Non-standard parallel solution strategies for distributed sparse linear systems. Lecture Notes in Computer Science, 1557:13–27, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-
ny.com/link/service/series/
0558/bibs/1557/15570013.htm;
 http://link.springer-ny.com/
link/service/series/0558/papers/
1557/15570013.pdf.

Saad:2002:PPS

[SS02a] Y. Saad and M. Sosonkina. pARMS: A package for solving general sparse linear systems on parallel computers. Lecture Notes in Computer Science, 2328:446–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-
ny.com/link/service/series/
0558/bibs/2328/23280446.htm;
 http://link.springer-ny.com/
link/service/series/0558/papers/
2328/23280446.pdf.

Saad:2002:AAR

[SS02b] Y. Saad and B. Suchomel. ARMS: an algebraic recursive multilevel solver for general sparse linear systems. Numerical Linear Algebra with Applications, 9(5):359–378, July/August 2002. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Sidje:2011:RAF

[SS11] Roger B. Sidje and Yousef Saad. Rational approximation to the Fermi–Dirac function with applications in density functional theory. Numerical Algorithms, 56 (3):455–479, March 2011. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=1017-1398&
volume=56&issue=3&spage=455.

Seghouane:2014:PHD

[SS14] Abd-Krim Seghouane and Yousef Saad. Prewhitening highdimensional fMRI data sets without eigendecomposition. Neural Comput., 26(5):907–919, 2014. ISSN 0899-7667 (print), 1530-888x (electronic).

Saad:1996:SLE

[SSC^{+}96] Y. Saad, A. Stathopoulos, J. Chelikowsky, K. Wu, and S. Öğüt. Solution of large eigenvalue problems in electronic structure calculations. BIT Numerical Mathematics, 36(3):563– 578, September 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.
liu.se/BIT/contents/bit36.
html;   http://www.springerlink.
com/openurl.asp?genre=article&
issn=0006-3835&volume=36&issue=

3&spage=563. International Linear Algebra Year (Toulouse, 1995).

Sosonkina:2004:UPA

[SSC04] M. Sosonkina, Y. Saad, and X. Cai. Using the parallel algebraic recursive multilevel solver in modern physical applications. Future Generation Computer Systems, 20(3):489– 500, April 1, 2004. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Stathopoulos:1993:RPL

[SSF93] A. Stathopoulos, Y. Saad, and C. F. Fisher. Robust preconditioning of large, symmetric eigenvalue problems. Technical Report 93-093, Army High Performance Computing Research Center, Minneapolis, MN, USA, 1993. to appear Journal of Computational and Applied Mathematics.

Stathopoulos:1995:RPL

[SSF95] Andreas Stathopoulos, Yousef Saad, and Charlotte Froese Fischer. Robust preconditioning of large, sparse, symmetric eigenvalue problems. Journal of Computational and Applied Mathematics, 64(3):197–215, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Saad:1985:SED

[SSS85] Y. Saad, A. Sameh, and P. Saylor. Solving elliptic difference equations on a linear array of processors. SIAM Journal on Scientific and Statistical Computing, 6(4):1049–1063, October 1985. CODEN SIJCD4. ISSN 0196-5204.

Sheehan:2010:CET

[SSS10] Bernard N. Sheehan, Yousef Saad, and Roger B. Sidje. Computing exp(−τA)b with Laguerre polynomials. Electronic Transactions on Numerical Analysis, 37:147–165, 2010. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic).

URL http://etna.mcs.kent.
edu/vol.37.2010/pp147-165.
dir/pp147-165.pdf.

Saad:2004:VAR

[SST04] Yousef Saad, Azzeddine Soulaimani, and Ridha Touihri. Variations on algebraic recursive multilevel solvers (ARMS) for the solution of CFD problems. Applied Numerical Mathematics: Transactions of IMACS, 51(2–3): 305–327, November 2004. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Stathopoulos:1998:DTR

[SSW98] Andreas Stathopoulos, Yousef Saad, and Kesheng Wu. Dynamic thick restarting of the Davidson, and the implicitly restarted Arnoldi methods. SIAM Journal on Scientific Computing, 19 (1):227–245, January 1998. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/30416.

Special issue on iterative methods (Copper Mountain, CO, 1996).

Saad:1998:DDM

[SSZ98] Yousef Saad, Maria Sosonkina, and Jun Zhang. Domain decomposition and multi-level type techniques for general sparse linear systems. In Domain decomposition methods, 10 (Boulder, CO, 1997), volume 218 of Contemp. Math., pages 174–190. American Mathematical Society, Providence, RI, USA, 1998.

Saad:2000:ISL

[Sv00] Yousef Saad and Henk A. van der Vorst. Iterative solution of linear systems in the 20th century. Journal of Computational and Applied Mathematics, 123(1–2): 1–33, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). Numerical analysis 2000, Vol. III. Linear algebra.

Saad:1988:BPSb

[SW88a] Y. Saad and H. Wijshoff. A benchmark package for sparse matrix computations. In J. Lenfant and D. De groot, editors, Proceedings of ICS conference 1988, St Malo, France, pages 500–509. ACM, 1988.

Saad:1988:BPSa

[SW88b] Youcef Saad and Harry A. G. Wijshoff. Benchmark package for sparse matrix computations. Technical Report CSRD 787, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 10 pp.

Saad:1989:SSH

[SW89] Y. Saad and H. Wijshoff. Some statistics on the Harwell–Boeing sparse matrices. Technical report, University of Illinois, CSRD, Urbana, IL, 1989. In preparation.

Saad:1990:SBP

[SW90] Youcef Saad and Harry A. G.

Wijshoff. SPARK: a benchmark package for sparse computations. ACM SIGARCH Computer Architecture News, 18(3b): 239–253, September 1990. CO-DEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Saad:1993:DQM

[SW93] Y. Saad and K. Wu. DQGM-RES: a quasi-minimal residual algorithm based on incomplete orthogonalization. Technical Report UMSI-93/131, Minnesota Supercomputing Institute, Minneapolis, MN, USA, 1993. submitted.

Saad:1994:PSM

[SW94] Y. Saad and K. Wu. Parallel sparse matrix library (P SPARSLIB): The iterative solvers module. Technical Report 94-008, Army High Performance Computing Research Center, Minneapolis, MN, USA, 1994. A short version to appear in Proceedings of PCG-94, Yokohama, Japan, March 14-17, 1994.

Saad:1995:DIS

[SW95] Yousef Saad and Kesheng Wu.

Design of an iterative solution module for a parallel sparse matrix library (P SPARSLIB). Applied Numerical Mathematics: Transactions of IMACS, 19(3): 343–357, 1995. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Special issue on iterative methods for linear equations (Atlanta, GA, 1994).

Saad:1996:DIS

[SW96a] Yousef Saad and Kesheng Wu. Design of an iterative solution module for a parallel sparse matrix library (P SPARSLIB). Applied Numerical Mathematics: Transactions of IMACS, 19(3): 343–357, January 15, 1996. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

URL http://www.elsevier.
com/cgi-bin/cas/tree/store/
apnum/cas_sub/browse/browse.
cgi?year=1996&volume=19&issue=

3&aid=638. Special issue on iterative methods for linear equations (Atlanta, GA, 1994).

Saad:1996:DDQ

[SW96b] Yousef Saad and Kesheng Wu. DQGMRES: a direct quasiminimal residual algorithm based on incomplete orthogonalization. Numerical Linear Algebra with Applications, 3(4):329– 343, July/August 1996. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (elec-

tronic). URL http://www3.
interscience.wiley.com/cgi-
bin/abstract?ID=15000994.

Saad:2000:DVC

[SYEG00] Y. Saad, M. Yeung, J. Erhel, and F. Guyomarc’h. A deflated version of the conjugate gradient algorithm. SIAM Journal on Scientific Computing, 21(5):1909–1926, September 2000. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/

article/33976. Iterative methods for solving systems of algebraic equations (Copper Mountain, CO, 1998).

Saad:1999:BBV

[SZ99a] Yousef Saad and Jun Zhang.

BILUM: Block versions of multielimination and multilevel ILU preconditioner for general sparse linear systems. SIAM Journal on Scientific Computing, 20(6): 2103–2121, November 1999. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/32753.

Saad:1999:BDB

[SZ99b] Yousef Saad and Jun Zhang. BI-LUTM: A domain-based multilevel block ILUT preconditioner for general sparse matrices. SIAM Journal on Matrix Analysis and Applications, 21 (1):279–299, 1999. CODEN SJ-MAEL. ISSN 0895-4798 (print), 1095-7162 (electronic). URL

http://epubs.siam.org/sam-
bin/dbq/article/34126.

Saad:1999:DTT

[SZ99c] Yousef Saad and Jun Zhang. Diagonal threshold techniques in robust multi-level ILU preconditioners for general sparse linear systems. Numerical Linear Algebra with Applications, 6 (4):257–280, June 1999. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (elec-

                                        tronic). URL http://www3.
                                        interscience.wiley.com/cgi-
                                        bin/abstract/65500095/START;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=65500095&PLACEBO=IE.pdf.

Saad:2001:EML

[SZ01] Yousef Saad and Jun Zhang. Enhanced multi-level block ILU preconditioning strategies for general sparse linear systems. Journal of Computational and Applied Mathematics, 130(1–2):99– 118, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Tang:2011:DDT

[TS11] Jok M. Tang and Yousef Saad. Domain-decomposition-type methods for computing the diagonal of a matrix inverse. SIAM Journal on Scientific Computing, 33(5):2823– 2847, ???? 2011. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

                                        URL http://epubs.siam.org/
                                        sisc/resource/1/sjoce3/v33/
                                        i5/p2823_s1.

Tang:2012:PMC

[TS12] Jok M. Tang and Yousef Saad. A probing method for computing the diagonal of a matrix inverse. Numerical Linear Algebra with Applications, 19(3):485–501, May 2012. CODEN NLAAEM.

ISSN 1070-5325 (print), 1099-1506 (electronic).

Tiago:2006:EMI

[TZA^{+}06] Murilo L. Tiago, Yunkai Zhou, M. M. G. Alemany, Yousef Saad, and James R. Chelikowsky. Evolution of magnetism in iron from the atom to the bulk. Physical Review Letters, 97(14):147201:1– 147201:4, October 6, 2006. CO-DEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http:

//link.aps.org/doi/10.1103/
PhysRevLett.97.147201.

Ubaru:2017:FES

[UCS17] Shashanka Ubaru, Jie Chen, and Yousef Saad. Fast estimation of tr(f(A)) via stochastic Lanczos quadrature. SIAM Journal on Matrix Analysis and Applications, 38(4):1075–1099, ???? 2017. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Ubaru:2017:LRA

[UMS17] Shashanka Ubaru, Arya Mazumdar, and Yousef Saad. Low rank approximation and decomposition of large matrices using error correcting codes. IEEE Transactions on Information Theory, 63(9):5544–5558, 2017. CO-DEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Ubaru:2019:SMC

[US19] Shashanka Ubaru and Yousef Saad. Sampling and multilevel coarsening algorithms for fast matrix approximations. Numerical Linear Algebra with Applications, 26(3):e2234:1–e2234:??, May 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ubaru:2017:FEA

[USS17a] Shashanka Ubaru, Yousef Saad, and Abd-Krim Seghouane. Fast estimation of approximate matrix ranks using spectral densities. Neural Computation, 29(5): 1317–1351, 2017. ISSN 0899-7667 (print), 1530-888x (electronic).

Ubaru:2017:IIL

[USS17b] Shashanka Ubaru, Abd-Krim Seghouane, and Yousef Saad. Improving the incoherence of a learned dictionary via rank shrinkage. Neural Computation, 29(1):263–285, 2017. ISSN 0899-7667 (print), 1530-888x (electronic).

Vecharynski:2014:FUA

[VS14] Eugene Vecharynski and Yousef Saad. Fast updating algorithms for latent semantic indexing. SIAM Journal on Matrix Analysis and Applications, 35(3): 1105–1131, 2014. CODEN SJ-MAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Vecharynski:2014:GPU

[VSS14] Eugene Vecharynski, Yousef Saad, and Masha Sosonkina. Graph partitioning using matrix values for preconditioning symmetric positive definite systems. SIAM Journal on Scientific Computing, 36(1):A63– A87, 2014. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Wang:2018:PRR

[WGSC18] Gang Wang, Georgios B. Giannakis, Yousef Saad, and Jie Chen. Phase retrieval via reweighted amplitude flow. IEEE Transactions on Signal Processing, 66(11):2818–2833, 2018. CO-DEN ITPRED. ISSN 1053-587x (print), 1941-0476 (electronic).

Wu:1993:PCM

[WS93] Kesheng Wu and Youcef Saad.

Performance of the CM-5 message passing primitives. Technical Report 20, Department of Computer Science, University of Minnesota, Minneapolis, MN 55455, USA, 1993.

Wu:1998:INP

[WSS98] Kesheng Wu, Yousef Saad, and Andreas Stathopoulos. Inexact Newton preconditioning techniques for large symmetric eigenvalue problems. Electronic Transactions on Numerical Analysis, 7:202–214, 1998. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (elec-

tronic). URL http://etna.mcs.
kent.edu/vol.7.1998/pp202-
214.dir/pp202-214.pdf. Large

scale eigenvalue problems (Argonne, IL, 1997).

Xu:2023:PAC

[XAKS23] Tianshi Xu, Anthony Austin, Vasileios Kalantzis, and Yousef Saad. A parallel algorithm for computing partial spectral factorizations of matrix pencils via Chebyshev approximation. SIAM Journal on Scientific Computing, 46(2):S324– S351, September 2023. CODEN SJOCE3. ISSN 1095-7197.

Xu:2024:PAC

[XAKS24] Tianshi Xu, Anthony Austin, Vasileios Kalantzis, and Yousef Saad. A parallel algorithm for computing partial spectral factorizations of matrix pencils via Chebyshev approximation. SIAM Journal on Scientific Computing, 46(2): S324–S351, 2024. CODEN SJOCE3. ISSN 1064-8275,1095-

7197. URL https://doi.org/
10.1137/22M1501155.

Xu:2022:PPM

[XKL^{+}22] Tianshi Xu, Vassilis Kalantzis, Ruipeng Li, Yuanzhe Xi, Geoffrey Dillon, and Yousef Saad. parGeMSLR: a parallel multilevel Schur complement lowrank preconditioning and solution package for general sparse matrices. Parallel Computing, 113(??):??, October 2022. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0167819122000497.

Xi:2016:AMP

[XLS16] Yuanzhe Xi, Ruipeng Li, and Yousef Saad. An algebraic multilevel preconditioner with lowrank corrections for sparse symmetric matrices. SIAM Journal on Matrix Analysis and Applications, 37(1):235–259, 2016. CO-DEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Xi:2018:FCS

[XLS18] Yuanzhe Xi, Ruipeng Li, and Yousef Saad. Fast computation of spectral densities for generalized eigenvalue problems. SIAM Journal on Scientific Computing, 40(4):A2749–A2773, ???? 2018. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Xi:2016:CPS

[XS16] Yuanzhe Xi and Yousef Saad.

Computing partial spectra with least-squares rational filters. SIAM Journal on Scientific Computing, 38(5):A3020–A3045, ???? 2016. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Xi:2017:RFP

[XS17] Yuanzhe Xi and Yousef Saad.

A rational function preconditioner for indefinite sparse linear systems. SIAM Journal on Scientific Computing, 39(3): A1145–A1167, ???? 2017. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Ye:2021:PGP

[YXS21] Xin Ye, Yuanzhe Xi, and Yousef Saad. Proxy-GMRES: Preconditioning via GMRES in polynomial space. SIAM Journal on Matrix Analysis and Applications, 42(3):1248–1267, ???? 2021. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Zhou:2014:CFS

[ZCS14] Yunkai Zhou, James R. Chelikowsky, and Yousef Saad. Chebyshev-filtered subspace iteration method free of sparse diagonalization for solving the Kohn–Sham equation. Journal of Computational Physics, 274 (??):770–782, October 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0021999114004744.

Zhou:2007:CDA

[ZS07] Yunkai Zhou and Yousef Saad. A Chebyshev–Davidson algorithm for large symmetric eigenproblems. SIAM Journal on Matrix Analysis and Applications, 29(3): 954–971, ???? 2007. CODEN SJ-MAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Zhou:2008:BKS

[ZS08] Yunkai Zhou and Yousef Saad. Block Krylov–Schur method for large symmetric eigenvalue problems. Numerical Algorithms, 47 (4):341–359, April 2008. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=1017-1398&
volume=47&issue=4&spage=341.

Zhou:2006:PSC

[ZSTC06a] Yunkai Zhou, Yousef Saad, Murilo L. Tiago, and James R. Chelikowsky. Parallel selfconsistent-field calculations via Chebyshev-filtered subspace acceleration. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 74(6): 066704:1–066704:8, December 2006. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic). URL http:/

/link.aps.org/doi/10.1103/
PhysRevE.74.066704.

Zhou:2006:SCF

[ZSTC06b] Yunkai Zhou, Yousef Saad, Murilo L. Tiago, and James R. Chelikowsky. Self-consistent-field calculation using Chebyshevfiltered subspace iteration. Journal of Computational Physics, 219(1):172–184, November 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S002199910600146X.

Zheng:2020:MLR

[ZXS20] Qingqing Zheng, Yuanzhe Xi, and Yousef Saad. Multicolor lowrank preconditioner for general sparse linear systems. Numerical Linear Algebra with Applications, 27(4):e2316:1–e2316:??, August 2020. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Zheng:2021:PSC

[ZXS21] Qingqing Zheng, Yuanzhe Xi, and Yousef Saad. A power Schur complement low-rank correction preconditioner for general sparse linear systems. SIAM Journal on Matrix Analysis and Applications, 42(2):659–682, 2021. CO-DEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).